### v2

Abstract
Segre–Veronese manifolds are smooth submanifolds of tensors comprising the partially symmetric rank-$1$ tensors. We investigate a one-parameter family of warped geometries of Segre–Veronese manifolds, which includes the standard Euclidean geometry. This parameter controls by how much spherical tangent directions are weighted relative to radial tangent directions. We present closed expressions for the exponential map, the logarithmic map, and the intrinsic distance on these warped Segre–Veronese manifolds, which can be computed efficiently numerically. It is shown that Segre–Veronese manifolds are not geodesically connected in the Euclidean geometry, while they are for some values of the warping parameter. The benefits of connecting geodesics may outweigh using the Euclidean geometry in certain applications. One such application is presented: numerically computing the Riemannian center of mass for averaging rank-$1$ tensors.
keywords:
Segre–Veronese manifold; partially symmetric rank-$1$ tensors; geodesic; exponential map; logarithmic map; warped geometry
{AMS}
15A69; 65F99; 53C22; 53A35; 14N07
1 Introduction
The $\mathbf{k}$th Segre–Veronese manifold $\mathcal{S}^{\mathbf{k}}$ of partially symmetric rank-$1$ tensors in the space of $n_{1}^{k_{1}}\times n_{2}^{k_{2}}\times\cdots\times n_{d}^{k_{d}}$ real arrays is a fundamental set of elementary tensors [52]. Special cases include the punctured plane
$\mathcal{S}^{(1)}:=\mathbb{R}^{n}_{*}:=\set{\lambda\mathbf{u}\mid\lambda\in{\mathbb{R}_{*}},\norm{\mathbf{u}}=1}$;
the manifold of symmetric rank-$1$ matrices
$\mathcal{S}^{(2)}:=\set{\lambda\mathbf{u}\mathbf{u}^{\mathsf{T}}\mid\lambda\in{\mathbb{R}_{*}},\,\norm{\mathbf{u}}=1}\subset\mathbb{R}^{n\times n},$
and, more generally, the Veronese manifold $\mathcal{S}^{(d)}$ of symmetric rank-$1$ tensors [52, 23]; and the manifold of rank-$1$ matrices
$\mathcal{S}^{(1,1)}:=\set{\lambda\mathbf{u}\mathbf{v}^{\mathsf{T}}\mid\lambda\in{\mathbb{R}_{*}},\,\norm{\mathbf{u}}=1,\,\norm{\mathbf{v}}=1}\subset\mathbb{R}^{m\times n},$
and, more generally, the Segre manifold $\mathcal{S}^{(1,\ldots,1)}$ of rank-$1$ tensors [52]. Analogously to the role of rank-$1$ matrices in a low-rank matrix decomposition, Segre–Veronese manifolds are the basic building blocks of more advanced tensor rank decomposition models [52]. Such sums of partially symmetric rank-$1$ tensors feature in a plethora of applications in biomedical engineering, chemometrics, computer science, data science, machine learning, psychometrics, signal processing, and statistics, among others; see the overview articles [51, 76, 70, 4, 63, 9] and the references therein.
The algebraic geometry of the Segre–Veronese manifold and its secant varieties have been thoroughly studied since the $19$th century [8]. By contrast, comparatively little is known about the Riemannian geometry of Segre–Veronese manifolds, let alone of (the smooth loci of) its higher secant varieties. To our knowledge, only the Euclidean geometry of the Segre manifold of rank-$1$ tensors (i.e., $\mathbf{k}=(1,\dots,1)$) as a submanifold of $\mathbb{R}^{n_{1}\times\dots\times n_{d}}$ was investigated in the prior works [79, 78].
The present paper studies essential properties of the Segre–Veronese manifold $\mathcal{S}^{\mathbf{k}}$ as a metric space. We are primarily interested in characterizing its geodesics. These curves $\gamma\subset\mathcal{S}^{\mathbf{k}}$ can be thought of as (locally) length-minimizing curves that generalize straight lines in Euclidean spaces. We distinguish between two types:
  1. 1.
    Connecting geodesics are related to the manifold’s logarithmic map. They are characterized by two endpoints $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}^{\mathbf{k}}$ and satisfy $\gamma(0)=\mathpzc{P}$ and $\gamma(\ell)=\mathpzc{Q}$, where $\ell$ is the distance between $\mathpzc{P}$ and $\mathpzc{Q}$. Such curves connect $\mathpzc{P}$ and $\mathpzc{Q}$ at the endpoints of $\gamma$.
  2. 2.
[…]
1.1 Applications of geodesics
Connecting geodesics find application in (i) statistics on manifolds [30], (ii) interpolation on manifolds [85], and (iii) approximation of maps into manifolds. First, the definition and computation of statistics on manifolds often involves connecting geodesics [30]. The simplest statistic, the Fréchet mean [31], Riemannian center of mass, or Karcher mean [46], on a manifold $\mathcal{M}$ consists of the point(s) that minimize the integral of the squared intrinsic distances to a fixed set of points on $\mathcal{M}$. The gradient of the squared intrinsic distance function from $x$ to a fixed point $x^{\star}\in\mathcal{M}$ is the tangent vector at $x$ of a connecting geodesic between $x$ and $x^{\star}$. Therefore, connecting geodesics appear for computing such a Fréchet mean both in gradient-based Riemannian optimization algorithms  [72, 71, 45] and other approximation algorithms [46, 45, 62, 19]. Connecting geodesics also appear in the definition and computation of higher-order statistics on manifolds, such as principal geodesic analysis [27, 28, 77] and the Gaussian distribution and its parameter estimation [71, 18]. They are an essential ingredient in geodesic regression [11, 29, 30] and generalizations thereof [40]. Second, many algorithms for interpolation of data on manifolds rely on connecting geodesics [68, 60, 44, 73, 67, 32, 74, 84, 85, 66, 87, 86]. Third, connecting geodesics are an essential building block of algorithms for constructing an approximation of a function from a Euclidean space into a Riemannian manifold. They are used in the Riemannian moving least squares algorithm [34, 35, 75], and the pullback-based methods of [57, 43, 83].
Starting geodesics have applications in (i) Riemannian optimization [1, 10], and (ii) integration on manifolds [37]. First, starting geodesics (or approximations thereof called retractions [80]) are an essential component of Riemannian optimization algorithms, which optimize a smooth objective function over a constraint set that is a smooth manifold [1, 10]. Second, starting geodesics, the projection retraction [2], or other retractions are essential in intrinsic integration schemes for vector fields on manifolds, such as the ones arising from differential equations on manifolds [24, 39, 65, 64, 38, 55, 37] or the projected ones from dynamic low-rank approximation methods (see [49, 50, 59, 82, 58, 26, 15, 41, 20]).
The foregoing discussion described general potential applications of geodesics. Next, we highlight plausible applications in the context of Segre–Veronese manifolds. Starting geodesics can be used in Riemannian optimization for seeking an approximate partially symmetric tensor rank decomposition [8]. Such algorithms were already applied to approximate tensor rank decompositions using a retraction in [13] and using geodesics in [79], and to symmetric tensor rank or Waring decompositions [8, 23] using a retraction in [48]. As illustrated in Section 6, computing a Fréchet mean is a natural alternative to averaging tensor rank decompositions based on factor matrices as was proposed in [16] to improve the robustness of stochastic tensor decompositions.
1.2 Contributions
Our main contributions are analytic expressions of the exponential and logarithmic maps of Segre–Veronese manifolds, in Corollary 3.4 and Theorem 4.15 respectively. We establish a few auxiliary results, such as a neat formula for the distance between points on Segre–Veronese manifolds in Proposition 4.18.
Proposition 4.17 entails that not every pair of partially symmetric rank-$1$ tensors on a Segre–Veronese manifold can be connected by a minimizing geodesic in the standard Euclidean geometry. In some applications, such as the one in Section 6, it is more important that every pair of points is connected by a minimizing geodesic than that Euclidean geometry is used. For this reason, we investigate $\alpha$-warped geometries [21] of Segre–Veronese manifolds. These geometries are described in detail in Section 2; essentially, they scale, by a factor of $\alpha>0$, the length of all tangent directions that do not infinitesimally change the Euclidean norm. The choice $\alpha=1$ corresponds to the Euclidean geometry. Figure 1 illustrates the minimizing geodesics connecting $(0,1)$ and $(1,0)$ for various values of $0<\alpha<2$ for the punctured plane $\mathbb{R}^{n}_{*}=\mathcal{S}^{(1)}$.
Figure 1: Geodesics between $(0,1)$ and $(1,0)$ in the $\alpha$-warped geometry for $\alpha=0.01$ (the red path approaching the semicircle), $\frac{1}{4}$, $\frac{1}{2}$, $\frac{3}{4}$, 11 (the purple straight line), $\frac{5}{4}$, $\frac{3}{2}$, $\frac{7}{4}$, and $1.99$ (the blue path that narrowly avoids the puncture at the origin).
1.3 Outline
The next section introduces background material from Riemannian geometry and describes $\alpha$-warped geometries of $\mathbb{R}_{*}^{n}$. General $\alpha$-warped Segre–Veronese manifolds are obtained as Riemannian embedded submanifolds thereof. We also give an alternative characterization of $\alpha$-warped Segre–Veronese manifolds, revealing them to be covered by a warped product of the positive real numbers and a product of spheres, which we call a pre-Segre–Veronese manifold. This characterization is exploited in the subsequent sections: the exponential maps of (pre-)Segre–Veronese manifolds are computed in Section 3 and their logarithmic maps are computed in Section 4. We present the sectional curvatures of Segre–Veronese manifolds in Section 5. Thereafter, in Section 6, we detail an example application of connecting geodesics for computing the Fréchet mean.
Notation
For convenience, scalars will be typeset in a lower case, vectors in a bold lower case, tensors in a calligraphic upper case, and manifolds in another calligraphic upper case. We will emphasize that a vector lives in a tangent space by adding an upper dot, e.g., $\dot{\mathbf{v}}$.
Unless otherwise specified, $(\mathbb{R}^{n},\langle\cdot,\cdot\rangle)$ will be abbreviated to $\mathbb{R}^{n}$; it is the inner product space of $n$-dimensional real vectors equipped with the standard Euclidean inner product. If $\mathcal{M}\subset\mathbb{R}^{n}$ is any subset, then $\mathbb{S}(\mathcal{M}):=\set{\mathbf{u}\in\mathcal{M}\mid\norm{\mathbf{u}}=1}$ is the space of vectors of $\mathcal{M}$ of unit Euclidean norm. We use the shorthand $\mathbb{S}^{n-1}:=\mathbb{S}(\mathbb{R}^{n})$ for the unit sphere in $\mathbb{R}^{n}$. The angular distance from $\mathbf{u}$ to $\mathbf{v}$, i.e., the arclength of a shortest great circle segment connecting them, is $\sphericalangle(\mathbf{u},\mathbf{v})=\cos^{-1}(\langle\mathbf{v},\mathbf{u}\rangle)\in[0,\pi]$.
Let $\mathcal{M}$ be a manifold. Its tangent space at $x\in\mathcal{M}$ is denoted by $\mathrm{T}_{x}\mathcal{M}$. The exponential map of $\mathcal{M}$ at $x$ is $\exp_{x}\colon\Omega_{x}\to\mathcal{M}$, where $\Omega_{x}$ is a suitable open neighborhood of $0\in\mathrm{T}_{x}\mathcal{M}$. The logarithmic map of $\mathcal{M}$ at $x$ is $\log_{x}\colon\mathcal{M}_{x}\to\mathrm{T}_{x}\mathcal{M}$, where $\mathcal{M}_{x}$ is a suitable open neighborhood of $x\in\mathcal{M}$. For a curve $\gamma\colon\mathbb{R}\to\mathcal{M}$, we denote the derivative to its parameter (see [54, Section 3] for a definition) using a prime, i.e.,
$\gamma^{\prime}(t):=\frac{\mathrm{d}}{\mathrm{d}t}\gamma(t).$
2 $\alpha$-warped geometries for Segre–Veronese manifolds
The next subsection recalls preliminary results from Riemannian geometry. In particular, the warped product of manifolds will be described. Then, in Section 2.2, the $\mathbf{k}$th Segre–Veronese manifold is formally defined and presented as a Riemannian embedded submanifold of a warped geometry of the space of tensors. Section 2.3 shows how Segre–Veronese manifolds can be realized as a normal Riemannian covering of a simpler warped product involving only spheres and the positive reals.
2.1 Warped geometries of $\mathbb{R}^{n}_{*}$
Recall that a smooth embedded manifold $\mathcal{M}\subset\mathbb{R}^{N}$ equipped with a Riemannian metric—an inner product $g_{x}\colon\mathrm{T}_{x}\mathcal{M}\times\mathrm{T}_{x}\mathcal{M}\to\mathbb{R}$ on its tangent spaces $\mathrm{T}_{x}\mathcal{M}$ that satisfies some technical regularity conditions, see, e.g., [53]—induces an intrinsic metric structure on $\mathcal{M}$ as follows. A Riemannian metric induces a norm $\norm{\dot{\mathbf{x}}}_{x}:=\sqrt{g_{x}(\dot{\mathbf{x}},\dot{\mathbf{x}})}$ for all $\dot{\mathbf{x}}\in\mathrm{T}_{x}\mathcal{M}$, which can be used to measure the length $\ell$ of piecewise smooth curves $\gamma\colon[0,1]\to\mathcal{M}$:
[…]
Letting the set of all piecewise smooth curves be our length structure $L$, the manifold $\mathcal{M}$ becomes a length space [14]. The intrinsic metric is then
$\mathrm{dist}_{\mathcal{M}}(x,y):=\inf_{\begin{subarray}{c}\gamma\in L,\\ \gamma(0)=x,\,\gamma(1)=y\end{subarray}}\ell(\gamma).$
A diffeomorphism (a smooth bijection with smooth inverse) between manifolds that preserves the intrinsic metric is called an isometry.
A piecewise smooth curve $\gamma\in L$ connecting $x=\gamma(0)$ and $y=\gamma(1)$ such that the length of $\gamma$ equals the distance $\mathrm{dist}_{\mathcal{M}}(x,y)$ is called a minimizing geodesic [53]. That is, a minimizing geodesic is a shortest path in $L$ between $x$ and $y$ contained in $\mathcal{M}$. More generally, a geodesic is any piecewise smooth curve that satisfies the first-order optimality condition, i.e., is a critical point, of the variational problem $\min_{\gamma\in L}\ell(\gamma)$ subject to appropriate constraints (such as fixing two endpoints). Such curves are always locally length-minimizing [53, Theorem 6.12] in the sense that they are minimizing geodesics on a segment $\gamma|_{[0,t)}$ for some $t>0$. Note that isometries map geodesics to geodesics.
For each tangent vector $v\in\mathrm{T}_{x}\mathcal{M}$, there is a unique constant-speed geodesic satisfying $\gamma(0)=x$, $\gamma^{\prime}(0)=v$ [53, Theorem 4.10]. We call such a geodesic a starting geodesic. Starting geodesics thus define a map from (a domain in) $\mathrm{T}_{x}\mathcal{M}$ to $\mathcal{M}$ via $v\mapsto\gamma(1)$. We call this the exponential map. It is a diffeomorphism between a neighborhood of $0\in\mathrm{T}_{x}\mathcal{M}$ and a neighborhood of $x\in\mathcal{M}$ [53, Lemma 5.10]. Its inverse is called the logarithmic map, and the corresponding geodesics are called connecting geodesics because they are geodesics with a given start and end points.
We also mention two important concepts in Riemannian geometry: The Riemann tensor and the sectional curvature. The Riemann tensor on a Riemannian manifold $\mathcal{M}$ is a multilinear map $R\colon\mathrm{T}_{p}\mathcal{M}\times\mathrm{T}_{p}\mathcal{M}\times\mathrm{T}_{p}\mathcal{M}\to\mathrm{T}_{p}\mathcal{M}$ that measures the change in a tangent vector as it is parallel transported around a small loop. The sectional curvature is a map $K\colon\mathrm{T}_{p}\mathcal{M}\times\mathrm{T}_{p}\mathcal{M}\to\mathbb{R}$ that measures the angular defect of small triangles around $p$. See [53, Chapter 7] for precise definitions.
The punctured space $\mathbb{R}^{n}_{*}$ is an open submanifold of $\mathbb{R}^{n}$; hence, it is a smooth manifold. It can be equipped with the standard Euclidean metric $g$ at $\mathbf{a}\in\mathbb{R}^{n}_{*}$, so that $g_{\mathbf{a}}(\dot{\mathbf{x}},\dot{\mathbf{y}})=\langle\dot{\mathbf{x}},\dot{\mathbf{y}}\rangle=\dot{\mathbf{x}}^{\mathsf{T}}\dot{\mathbf{y}},$ where $\dot{\mathbf{x}}$, $\dot{\mathbf{y}}\in\mathrm{T}_{\mathbf{a}}\mathbb{R}^{n}_{*}=\mathbb{R}^{n}$ and $\langle\cdot,\cdot\rangle$ is the standard Euclidean inner product. This leads to the Euclidean geometry of $(\mathbb{R}^{n}_{*},g)$. Other meaningful geometries of $\mathbb{R}^{n}_{*}$ exist. For example, $\mathbb{R}^{n}_{*}$ is diffeomorphic to the product of the positive real line and an (n1)(n-1)-dimensional sphere: $\mathbb{R}^{n}_{*}\simeq\mathbb{R}_{>0}\times\mathbb{S}^{n-1}.$ Under this identification, the above standard Euclidean metric can be expressed as
$g_{\lambda\mathbf{u}}(\dot{x}\mathbf{u}+\lambda\dot{\mathbf{v}},\dot{y}\mathbf{u}+\lambda\dot{\mathbf{w}})=\dot{x}\dot{y}+\lambda^{2}\langle\dot{\mathbf{v}},\dot{\mathbf{w}}\rangle,$
where $\dot{x}$, $\dot{y}\in\mathbb{R}=\mathrm{T}_{\lambda}\mathbb{R}_{>0}$ and $\dot{\mathbf{v}}$, $\dot{\mathbf{w}}\in\mathbf{u}^{\perp}=\mathrm{T}_{\mathbf{u}}\mathbb{S}^{n-1}$. Herein, the inner products on the right are the inner products inherited from the ambient Euclidean spaces $\mathbb{R}$ and $\mathbb{R}^{n}$, respectively. Note in particular that the metric that $\mathbb{R}^{n}_{*}$ inherits from $\mathbb{R}^{n}$ is not equal to the product metric, $\dot{x}\dot{y}+\left\langle{}\dot{\mathbf{v}},\dot{\mathbf{w}}\right\rangle$, of $\mathbb{R}_{>0}$ and $\mathbb{S}^{n-1}$. Instead, it is a so-called warped product [69, Chapter 7] with the identity map as warping function. Recall from [69, Chapter 7] that if we are given Riemannian manifolds $(\mathcal{M},g)$ and $(\mathcal{N},h)$, then we can construct their warped product with warping function $f\colon\mathcal{M}\to\mathbb{R}$ as $\mathcal{M}\times_{f}\mathcal{N}:=(\mathcal{M}\times\mathcal{N},g\times_{f}h),$ where
$(g\times_{f}h)_{(m,n)}((\dot{\mathbf{m}}_{1},\dot{\mathbf{n}}_{1}),(\dot{\mathbf{m}}_{2},\dot{\mathbf{n}}_{2}))=g(\dot{\mathbf{m}}_{1},\dot{\mathbf{m}}_{2})+f^{2}(m)\cdot h(\dot{\mathbf{n}}_{1},\dot{\mathbf{n}}_{2}).$
Rather than considering only the Euclidean geometry of $\mathbb{R}^{n}_{*}$, in this paper, we also consider its $\alpha$-warped geometries; that is,
(1) $\mathcal{R}_{\alpha}=(\mathbb{R}^{n}_{*},\varsigma^{\alpha}):=\mathbb{R}_{>0}\times_{\alpha\mathrm{Id}}\mathbb{S}^{n-1}\quad\text{with }\alpha>0,$
whose warped product metric satisfies
(2) $\varsigma^{\alpha}_{\lambda\mathbf{u}}(\dot{x}\mathbf{u}+\lambda\dot{\mathbf{v}},\dot{y}\mathbf{u}+\lambda\dot{\mathbf{w}})=\dot{x}\dot{y}+(\alpha\lambda)^{2}\langle\dot{\mathbf{v}},\dot{\mathbf{w}}\rangle.$
If $\alpha=1$, then the $1$-warped geometry is the Euclidean geometry of $\mathbb{R}^{n}_{*}$, with straight lines as geodesics between points. If $\alpha>1$, then distances in the spherical directions will be elongated, implying that geodesics between points will tend to be forced toward the origin. On the other hand, if $0<\alpha<1$, then spherical movements are easier, implying that geodesics between points will bend away from the origin. This is visualized in Fig. 1. The precise form of the geodesics is stated in Proposition 3.1.
2.2 The $\alpha$-warped Segre–Veronese submanifold
The punctured space $\mathbb{R}^{n}_{*}$ is the most basic example (with $\mathbf{k}=(1)$) of the $\mathbf{k}$th Segre–Veronese manifold
(3) $\mathcal{S}^{\mathbf{k}}_{\pm}:=\set{\lambda\mathbf{x}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{x}_{d}^{\otimes k_{d}}\mid\lambda\in\mathbb{R}_{*},\;\mathbf{x}_{i}\in\mathbb{S}^{n_{i}-1},\,i=1,\ldots,d}$,
where $\mathbf{x}_{i}^{\otimes k_{i}}$ is the $k_{i}$-fold tensor product of $\mathbf{x}_{i}$ with itself, i.e., $\mathbf{x}_{i}^{\otimes k_{i}}:=\mathbf{x}_{i}\otimes\cdots\otimes\mathbf{x}_{i}$. Recall that the tensor product of vectors in $\mathbb{R}^{n_{j}}$, $j=1$, …, $d$, can be embedded naturally into the space of $n_{1}\times\cdots\times n_{d}$ arrays, as follows:
$(\mathbf{x}_{1}\otimes\cdots\otimes\mathbf{x}_{d})(i_{1},\ldots,i_{d}):=\mathbf{x}_{1}(i_{1})\cdots\mathbf{x}_{d}(i_{d}),\qquad 1\leq i_{j}\leq n_{j},\,j=1,\ldots,d,$
where $\mathbf{x}_{j}\in\mathbb{R}^{n_{j}}$ and $\mathbf{x}_{j}(i_{j})$ denotes the $i_{j}$th coefficient of $\mathbf{x}_{j}$ with respect to a basis of $\mathbb{R}^{n_{j}}$; see, e.g., [56]. That is, the coordinates of a tensor product are obtained by multiplying the coordinates of all vectors in all possible ways.
Let us also define the ?positive? Segre–Veronese manifold
(4) $\mathcal{S}^{\mathbf{k}}:=\set{\lambda\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}}\mid\lambda\in\mathbb{R}_{>0},\;\mathbf{u}_{i}\in\mathbb{S}^{n_{i}-1},\,i=1,\ldots,d}$,
which is a smooth embedded submanifold (see e.g. [52, Section 4.3] and [54, Chapter 5]) that is diffeomorphic to $\mathbb{R}_{>0}\times\mathbb{S}(\mathcal{S}^{\mathbf{k}}).$ We can thus equip it with the $\alpha$-warped metric that it inherits as a submanifold of $\mathcal{R}_{\alpha}=\mathbb{R}_{>0}\times\mathbb{S}^{N-1}$ from Eq. 1, where $N=n_{1}^{k_{1}}\cdots n_{d}^{k_{d}}$. The resulting Riemannian submanifold will be denoted by
(5) $\mathcal{S}_{\alpha}^{\mathbf{k}}:=\mathbb{R}_{>0}\times_{\alpha\mathrm{Id}}\mathbb{S}(\mathcal{S}^{\mathbf{k}})=(\mathcal{S}^{\mathbf{k}},\varsigma^{\alpha})\subset\mathcal{R}_{\alpha}$
and we will refer to it as the $\alpha$-warped Segre–Veronese manifold.
Remark 2.1.
If all $k_{i}$’s are even, then $\mathcal{S}_{\pm}^{\mathbf{k}}$ consists of two connected components, namely the positive Segre–Veronese manifold $\mathcal{S}^{\mathbf{k}}$ and its mirror image along the origin, namely the ?negative? Segre–Veronese manifold $\mathcal{S}^{\mathbf{k}}$. This means that $p\in\mathcal{S}^{\mathbf{k}}$ and $q\in-\mathcal{S}^{\mathbf{k}}$ are not connected by any curve in $\mathcal{S}^{\mathbf{k}}_{\pm}$. In particular, there will be no geodesics connecting them. Nevertheless, the two connected components are isometric, as $\varsigma_{(-\lambda)\mathbf{u}}^{\alpha}=\varsigma_{\lambda\mathbf{u}}^{\alpha}$ (herein, the metric should be interpreted as being defined on $\mathbb{R}_{\neq 0}$), and we have $\mathcal{S}^{\mathbf{k}}=\mathbb{R}_{<0}\times_{\alpha\mathrm{Id}}\mathbb{S}^{N-1}$. As these two connected components are isometric manifolds, their geometry is the same.
On the other hand, if at least one of the $k_{i}$’s is odd, then it is easy to show that $\mathcal{S}_{\pm}^{\mathbf{k}}=\mathcal{S}^{\mathbf{k}}=-\mathcal{S}^{k}$. That is, the ?regular?, ?positive?, and ?negative? flavors of the Segre–Veronese manifold all coincide.
Therefore, we will henceforth study the component $\mathcal{S}^{\mathbf{k}}$ without loss of generality.
2.3 The $\alpha$-warped pre-Segre–Veronese manifold
We study the geometry of $\mathcal{S}_{\alpha}^{\mathbf{k}}$ through the lens of the tensor product $\otimes$. By definition, $\mathcal{S}^{\mathbf{k}}$ is the image of
$\otimes\colon\mathbb{R}_{>0}\times\mathbb{S}^{n_{1}-1}\times\cdots\times\mathbb{S}^{n_{d}-1}$ $\longrightarrow\mathbb{R}^{n_{1}^{k_{1}}\times\cdots\times n_{d}^{k_{d}}}_{*},$
$(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ $\longmapsto\lambda\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}}.$
As the domain is a product manifold, its tangent space at $(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ is
$\mathrm{T}_{\lambda}\mathbb{R}_{>0}\times\mathrm{T}_{\mathbf{u}_{1}}\mathbb{S}^{n_{1}-1}\times\cdots\times\mathrm{T}_{\mathbf{u}_{d}}\mathbb{S}^{n_{d}-1}=\mathbb{R}\times\mathbf{u}_{1}^{\perp}\times\cdots\times\mathbf{u}_{d}^{\perp}.$
We can equip it with the $\alpha$-warped product metric
(6) $g_{(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})}\bigl{(}(\dot{x},\dot{\mathbf{u}}_{1},\ldots,\dot{\mathbf{u}}_{d}),(\dot{y},\dot{\mathbf{v}}_{1},\ldots,\dot{\mathbf{v}}_{d})\bigr{)}:=\dot{x}\dot{y}+(\alpha\lambda)^{2}\sum_{i=1}^{d}k_{i}\langle\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{i}\rangle.$
The resulting Riemannian manifold will be called the $\alpha$-warped pre-Segre–Veronese manifold, and it will be denoted by
(7) $\mathcal{P}_{\alpha}^{\mathbf{k}}=(\mathbb{R}_{>0},\langle\cdot,\cdot\rangle)\times_{\alpha\mathrm{Id}}\underbrace{((\mathbb{S}^{n_{1}-1},k_{1}\langle\cdot,\cdot\rangle)\times\cdots\times(\mathbb{S}^{n_{d}-1},k_{d}\langle\cdot,\cdot\rangle))}_{\mathbb{S}^{\mathbf{n},\mathbf{k}}}.$
Note that the standard Euclidean inner product on the $i$th sphere is weighted by a factor $k_{i}$. The geometry of such a sphere $(\mathbb{S}^{n_{i}-1},k_{i}\langle\cdot,\cdot\rangle)$ is equivalent to the usual Euclidean geometry of the sphere $(\mathbb{S}^{n_{i}-1}_{\sqrt{k_{i}}},\langle\cdot,\cdot\rangle)$ of radius $\sqrt{k_{i}}$.
The key result is that the tensor product takes the pre-Segre–Veronese manifold $\mathcal{P}_{\alpha}^{\mathbf{k}}$ in a nice way to the Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$. Specifically, it is a normal Riemannian covering map [69], which means that (i) every point $\mathpzc{T}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ has the same number, $N$, of preimages in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ (ii) every point $\mathpzc{T}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ has a neighborhood $U_{\mathpzc{T}}$ such that its preimage $\otimes^{-1}(U_{\mathpzc{T}})$ is a disjoint union of $N$ domains, and (iii) on each of those domains, $\otimes$ is an isometry. Property (iii) is called local isometry. A diffeomorphism $\iota\colon\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathcal{P}_{\alpha}^{\mathbf{k}}$ that satisfies $\otimes\circ\iota=\otimes$ is called a deck transformation. Morally, it shuffles the preimages of the covering map. A deck transform that is also an isometry is called an isometric deck transform.
Lemma 2.2.
The map
$\otimes\colon\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathcal{S}_{\alpha}^{\mathbf{k}},\quad(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\mapsto\lambda\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}}$
is a normal Riemannian covering where the isometric deck transforms are of the form $\imath_{\sigma}(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})=(\lambda,\sigma_{1}\mathbf{u}_{1},\ldots,\sigma_{d}\mathbf{u}_{d})$ with $\sigma_{i}\in\set{-1,1}$ and $\sigma_{1}^{k_{1}}\cdots\sigma_{d}^{k_{d}}=1$.
An elementary but crucial fact about normal Riemannian coverings concerns the lengths of curves in the covered manifold. The next result enables the computation of geodesics in $\mathcal{S}_{\alpha}^{\mathbf{k}}$ by studying the geodesics of the warped product $\mathcal{P}_{\alpha}^{\mathbf{k}}$.
Lemma 2.3.
[…]
In particular, a minimizing geodesic segment on $\mathcal{N}$ lifts to a minimizing geodesic segment on $\mathcal{M}$.
These lemmas are proved in Appendix A.
By Lemma 2.3, if we want to find a minimizing connecting geodesic between $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$, it suffices to enumerate the connecting geodesics between the preimages and pick the shortest one. This is essentially what our proposed matchmaking algorithm will do in Proposition 4.19.
3 The exponential map
Before presenting the main original contribution in Section 4, we briefly investigate the exponential map on the (pre-)Segre–Veronese manifold. The following result is a straightforward generalization of [79, Theorem 1] and its proof, which dealt only with the usual Euclidean geometry (i.e., $\alpha=1$) of the Segre manifold (i.e., $\mathbf{k}=(1,\ldots,1)$). As we will rely on the general strategy of the proof, it is included for self-containedness.
Proposition 3.1 (Exponential map on $\mathcal{P}_{\alpha}^{\mathbf{k}}$).
Let
$p=(\lambda,\mathbf{u}_{1},\dots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}\quad\text{and}\quad\dot{p}=(\dot{\lambda},\dot{\mathbf{u}}_{1},\dots,\dot{\mathbf{u}}_{d})\in\mathrm{T}_{p}\mathcal{P}_{\alpha}^{\mathbf{k}}.$
Then
[…]
where
$a_{i}={}$ $\frac{\norm{\dot{\mathbf{u}}_{i}}}{\alpha N}\left(\frac{\pi}{2}-\arctan\left(\frac{\lambda+\dot{\lambda}}{\lambda\alpha N}\right)\right),\quad i=1,\dots,d,$
$N={}$ $\sqrt{\sum_{i=1}^{d}k_{i}\norm{\dot{\mathbf{u}}_{i}}^{2}}.$
If $N=0$ and $\lambda<\dot{\lambda}$, the geodesics are straight lines:
$\operatorname{exp}_{p}(\dot{p})=(\lambda+\dot{\lambda},\mathbf{u}_{1},\ldots,\mathbf{u}_{d}).$
Proof 3.2.
We are looking for the constant-speed geodesic $\gamma$ with initial conditions $\gamma(0)=p$, $\gamma^{\prime}(0)=\dot{p}$, so that we can evaluate it at $1$.
If $N=0$, the proof proceeds exactly as in [79, Proof of Theorem 1]. So we only need to consider the case $N>0$.
Let $\gamma(t)=(\lambda(t),\mathbf{u}_{1}(t),\dots,\mathbf{u}_{d}(t))$ be a geodesic in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ from Eq. 7 with starting conditions $\gamma(0)=p$ and $\gamma^{\prime}(0)=\dot{p}$. We apply the change of variable $t\mapsto s(t)$ with $s(0)=0$ and $s^{\prime}(0)>0$ such that
(8) $\sigma(s)=\left(\mu(s),\mathbf{v}_{1}(s),\dots,\mathbf{v}_{d}(s)\right),\;\text{with }\norm{(\mathbf{v}_{1}^{\prime}(s),\dots,\mathbf{v}_{d}^{\prime}(s))}_{(\mathbf{v}_{1}(s),\ldots,\mathbf{v}_{d}(s))}=1,$
i.e., $\sigma$ is the geodesic satisfying $\sigma(s(t))=\gamma(t)$ and which has unit spherical speed. Then, it holds that $\norm{\mathbf{v}_{i}^{\prime}(0)}_{\mathbf{v}_{i}(0)}s^{\prime}(0)=\sqrt{k_{i}}\norm{\dot{\mathbf{u}}_{i}}$, $i=1$, …, $d$. From this we find that $s^{\prime}(0)=N$ where $N$ is as in the statement of the proposition. Then, [69, Remark 7.39] implies that the projection of this geodesic on $\mathbb{S}^{n_{1}-1}\times\dots\times\mathbb{S}^{n_{d}-1}$ results in a geodesic. The geodesics on spheres are great circles:
(9) $\mathbf{v}_{i}(s)=\mathbf{u}_{i}\cos\left(\frac{\norm{\dot{\mathbf{u}}_{i}}}{N}s\right)+\frac{\dot{\mathbf{u}}_{i}}{\norm{\dot{\mathbf{u}}_{i}}}\sin\left(\frac{\norm{\dot{\mathbf{u}}_{i}}}{N}s\right),\quad i=1,\ldots,d.$
The coefficient $\mu\colon\mathbb{R}\to\mathbb{R}$ is determined by solving the Euler–Lagrange equations of the length functional, which in the $\alpha$-warped geometry becomes
[…]
This ultimately leads to the differential equation $\alpha^{2}\mu(s)^{2}+2\mu^{\prime}(s)^{2}-\mu(s)\mu^{\prime\prime}(s)=0$ whose solution is of the form
(11) $\mu(s)=\frac{A}{\cos(\alpha s+B)}.$
Solving for the constants $A$ and $B$ by exploiting the starting conditions, we arrive at $A=\lambda\cos(B)$ and $B=\arctan(\dot{\lambda}/(\alpha\lambda N))$. In this way, we obtain the geodesic given by Eqs. 9 and 11.
The final step, consists of changing variables back to $t$. We have that
[…]
Since $\sigma(s)$ is injective, the reparametrization $s(t)$ is a monotone increasing function. Especially, $s^{\prime}(t)\geq 0$. By plugging $\norm{\gamma^{\prime}(t)}=\norm{\dot{p}}_{p}=\sqrt{\dot{\lambda}^{2}+(\alpha\lambda N)^{2}}$ into Eq. 12, we obtain the initial value problem
(13) $s(0)=0,\qquad s^{\prime}(t)=\cos^{2}(\alpha s(t)+B)\frac{\sqrt{\dot{\lambda}^{2}+(\alpha\lambda N)^{2}}}{\alpha A}.$
Combining this with the initial condition $s(0)=0$, we have a solution
(14) $s(t)=\alpha^{-1}\left(\frac{\pi}{2}-\arctan\left(\frac{\lambda+\dot{\lambda}t}{\alpha\lambda Nt}\right)\right).$
Verifying that this solves Eq. 13 is a straightforward computation. For completeness, we include it in Section A.4.
To obtain the exponential map, we now just have to evaluate $\gamma(1)$. For the first component of $\gamma(1)$ we find, after some computations (see Section A.5), that
(15) $\mu(s(1))={}$ $\sqrt{(\lambda+\dot{\lambda})^{2}+(\alpha\lambda N)^{2}}.$
The other components $\mathbf{v}_{i}(s(1))$ are obtained immediately by plugging the expression of $s(t)$ from Eq. 14 into Eq. 9. This concludes the proof.
From this result, it immediately follows that the exponential map is well defined almost everywhere.
Corollary 3.3.
The domain of the exponential map $\exp_{p}$ of $\mathcal{P}_{\alpha}^{\mathbf{k}}$ is the whole tangent space $\mathrm{T}_{p}\mathcal{P}_{\alpha}^{\mathbf{k}}$, except for the half-line $\set{(\dot{\lambda},0,\ldots,0)\mid\dot{\lambda}\leq-\lambda}$.
Having computed the geodesics on the pre-Segre–Veronese manifold, an application of Lemmas 2.2 and 2.3 enables us to push them forward with $\otimes$ to the Segre–Veronese manifold. Note that because Riemannian coverings are local isometries, we are only guaranteed that geodesics are mapped to geodesics, but not that minimizing geodesics are mapped to minimizing geodesics, i.e., the converse of Lemma 2.3 is not true. In this way, we arrive at the geodesics and the following characterization of the exponential map of $\mathcal{S}_{\alpha}^{\mathbf{k}}$.
Corollary 3.4 (The exponential map of $\mathcal{S}_{\alpha}^{\mathbf{k}}$).
Let $\mathpzc{P}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ be a rank-$1$ tensor and let $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\otimes^{-1}(\mathpzc{P})$ be any of its representatives on $\mathcal{P}_{\alpha}^{\mathbf{k}}$. Let $\dot{\mathpzc{P}}\in\mathrm{T}{P}\mathcal{S}_{\alpha}^{\mathbf{k}}$, which can be expressed uniquely as
$\dot{\mathpzc{P}}=\dot{\lambda}\mathpzc{U}+\lambda\left(\nu_{k_{1}}(\dot{\mathbf{u}}_{1})\otimes\mathbf{u}_{2}^{{\otimes}k_{2}}\cdots\otimes\mathbf{u}_{d}^{{\otimes}k_{d}}+\cdots+\mathbf{u}_{1}^{{\otimes}k_{1}}\otimes\cdots\otimes\mathbf{u}_{d-1}^{{\otimes}k_{d-1}}\otimes\nu_{k_{d}}(\dot{\mathbf{u}}_{d})\right),$
where
(16) $\mathpzc{U}$ $=\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}},$
(17) $\nu_{{k_{i}}}(\dot{\mathbf{u}}_{i})$ $=(\dot{\mathbf{u}}_{i}\otimes\mathbf{u}_{i}^{\otimes({k_{i}}-1)})+\cdots+(\mathbf{u}_{i}^{\otimes({k_{i}}-1)}\otimes\dot{\mathbf{u}}_{i})\quad\text{for all }\dot{\mathbf{u}}_{i}\in\mathrm{T}_{\mathbf{u}_{i}}\mathbb{S}^{n}.$
Let $\dot{p}=(\dot{\lambda},\dot{\mathbf{u}}_{1},\ldots,\dot{\mathbf{u}}_{d})\in\mathrm{T}_{p}\mathcal{P}_{\alpha}^{\mathbf{k}}$ be the corresponding lifted tangent vector. Then,
$\exp{P}(\dot{\mathpzc{P}})=\otimes(\exp_{p}(\dot{p})),$
insofar as $\dot{p}$ is in the domain of $\exp_{p}$.
Another interesting consequence of Propositions 3.1 and 3.3 is that it entails the next result about the radius of the largest open ball in $\mathrm{T}{P}\mathcal{S}_{\alpha}^{\mathbf{k}}$ for which $\exp$ is a diffeomorphism, i.e., the local injectivity radius of the Segre–Veronese manifold.
Corollary 3.5 (Injectivity radius).
The local injectivity radius of $\mathcal{P}_{\alpha}^{\mathbf{k}}$, and thus also of $\mathcal{S}_{\alpha}^{\mathbf{k}}$, at $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ is $\lambda$. The (global) injectivity radius of $\mathcal{P}_{\alpha}^{\mathbf{k}}$ and $\mathcal{S}_{\alpha}^{\mathbf{k}}$, i.e., the infimum of the local injectivity radii over the manifold, is zero.
Note that the latter claim is also immediate from the fact that neither $\mathcal{P}_{\alpha}^{\mathbf{k}}$ nor $\mathcal{S}_{\alpha}^{\mathbf{k}}$ are complete metric spaces. Indeed, the sequence $p_{n}=(n^{-1},\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ tends to $(0,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\not\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ as $n\to\infty$, and the image of $p_{n}$ in $\mathcal{S}_{\alpha}^{\mathbf{k}}$ tends to $0\notin\mathcal{S}_{\alpha}^{\mathbf{k}}$.
Finally, we highlight that the exponential map on the pre-Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ can be computed numerically very efficiently by straightforward implementation of the formulas in Proposition 3.1. The cost is about $\mathcal{O}(\sum_{i=1}^{d}n_{i})$ operations to compute $N$ and all $\mathbf{u}_{i}(t)$’s, and a constant number of operations for $\lambda(t)$ and $g(t)$.
4 The logarithmic map
We now compute the logarithmic map of the $\alpha$-warped (pre-)Segre–Veronese manifold, significantly extending prior results of one of the authors (L. S.) on the Segre manifold ($\mathbf{k}=(1,\ldots,1)$) with the Euclidean metric ($\alpha=1$) as part of his PhD thesis [78].
In the next subsection, we first determine a practical characterization of when two points on the pre-Segre–Veronese manifold can be connected with a geodesic. Then, in Section 4.2, we compute this connecting minimizing geodesic if it exists. Finally, we compute minimizing geodesics on the Segre–Veronese manifold between two rank-$1$ tensors in Section 4.3 by appropriately lifting them to the pre-Segre–Veronese manifold. Additionally, an efficient matchmaking algorithm is presented to compute this lift, essentially solving a particular combinatorial problem in linear time.
4.1 Compatibility
Our main motivation for considering $\alpha$-warped geometries of the Segre–Veronese manifold is the observation in [78, Chapter 6] that not all pairs of points on the Segre manifold are connected by a minimizing geodesic. That is, the Segre manifold is not geodesically connected in the standard Euclidean geometry.
The main result of this subsection, Proposition 4.2, will show that the geodesic connectedness of $\mathcal{P}_{\alpha}^{\mathbf{k}}$ can be characterized by the following property.
Definition 4.1 ($\alpha$-compatibility).
Consider two points $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})$ in $\mathcal{P}_{\alpha}^{\mathbf{k}}$. Let
$M(p,q):=\mathrm{dist}_{\mathbb{S}^{\mathbf{n},\mathbf{k}}}((\mathbf{u}_{1},\ldots,\mathbf{u}_{d}),(\mathbf{v}_{1},\ldots,\mathbf{v}_{d}))=\sqrt{\sum_{i=1}^{d}k_{i}\cdot\sphericalangle^{2}(\mathbf{u}_{i},\mathbf{v}_{i})}.$
Then, $p$ and $q$ are called $\alpha$-compatible if $M(p,q)<\alpha^{-1}\pi.$
The dependency of $M(p,q)$ on the parameters $\mathbf{k}\geq 1$ will not be emphasized in the notation, instead being implicit in the fact that $p$, $q\in\mathcal{P}_{\alpha}^{\mathbf{k}}$.
Proposition 4.2.
Two points $p$ and $q$ can be connected by a minimizing geodesic in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ if and only if $p$ and $q$ are $\alpha$-compatible.
Proof 4.3.
The case where $\mathbf{u}_{i}=\mathbf{v}_{i}$ for all $i=1,\dots,d$ corresponds to the case of straight line geodesics (i.e., $M(p,q)=0$) in Proposition 3.1. All such $p$ and $q$ can be connected by a straight line and since $M(p,q)=0$ they are always $\alpha$-compatible. Hence, the claim holds in this special case.
The remainder of the proof treats the case $M(p,q)>0$. Let $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$. For brevity, we let $M=M(p,q)$. We prove the two directions separately.
Geodesics $\Longrightarrow$ compatibility
Assume that $p$ and $q$ are connected by a minimizing geodesic $\sigma$, parameterized to have unit spherical speed as in Eq. 8. Then, Chen’s result [21, Lemma 4.4] implies that the projection of $\sigma$ to the spherical part $\mathbb{S}^{\mathbf{n},\mathbf{k}}$, as in Eq. 7, is a minimizing unit-speed geodesic. As $\mathbb{S}^{\mathbf{n},\mathbf{k}}$ is a product of spheres, each equipped with a scaled Euclidean inner product, this minimizing geodesic is a product of arc segments of the great circles connecting $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$. The distance on $\mathbb{S}^{\mathbf{n},\mathbf{k}}$ between $(\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ and $(\mathbf{v}_{1},\ldots,\mathbf{v}_{d})$ thus satisfies $M^{2}=\sum_{i=1}^{d}k_{i}\cdot\sphericalangle^{2}(\mathbf{u}_{i},\mathbf{v}_{i})$.
In Section 2.1, we defined geodesics as critical points of a variational problem. Such a condition may equivalently be phrased in terms of an Euler–Lagrange equation such as Eq. 10. As $\sigma$ is a geodesic, the proof of Proposition 3.1 then implies that $\mu(s)$ is the form as in Eq. 11, where $A,B\in\mathbb{R}$ are determined by the conditions $\mu(0)=\lambda$ and $\mu(M)=\mu$. We have $A=\lambda\cos(B)$, as in the proof of Proposition 3.1, and
$\frac{\mu(0)}{\mu(M)}=\frac{\lambda}{\mu}=\frac{\cos(B+\alpha M)}{\cos(B)}=\cos(\alpha M)-\sin(\alpha M)\tan(B).$
Hence,
$B=\arctan\left(\cot(\alpha M)-\frac{\lambda\mu^{-1}}{\sin(\alpha M)}\right),$
provided that $\alpha M\neq\pi k$ for $k\in\mathbb{Z}$. Since $\sigma(s)$ exists by assumption, the solution $\mu(s)>0$ for all $0\leq s\leq M$. The inverse of the cosine has singularities at $B+\alpha s=\frac{\pi}{2}+k\pi$ for $k\in\mathbb{Z}$, so $\mu(s)$ cannot pass through them and we must have $\frac{\pi}{2}+k\pi<B+\alpha s<\frac{\pi}{2}+k\pi$ for all $0\leq s\leq M$. In particular, the endpoint $s=0$ implies that $\frac{\pi}{2}+k\pi<B<\frac{\pi}{2}+k\pi$. Consequently, $0<\alpha M<\pi$ because we also have that $M>0$ and $\alpha>0$. Hence, $p$ and $q$ are strictly $\alpha$-compatible.
Compatibility $\Longrightarrow$ geodesics
As before, consider minimizing geodesics $\mathbf{v}_{i}(s)$ connecting $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ on the sphere $(\mathbb{S}^{n_{i}-1},k_{i}\langle\cdot,\cdot\rangle)$ for $i=1,\ldots,d$. We can parameterize them so that they have unit spherical speed as in Eq. 8.
If there exists a geodesic connecting $p$ and $q$, it must be a solution to the Euler–Lagrange equation. Proceeding as above, a solution $\mu(s)$ exists only if $\alpha M\neq\pi k$ for $k\in\mathbb{Z}$. As $p$ and $q$ are $\alpha$-compatible, $0<\alpha M<\pi$, so $\mu(s)\subset\mathbb{R}$ is a valid smooth curve. For $\mu(s)$ to be a geodesic, however, it must also be contained in $\mathbb{R}_{>0}$ for all $0\leq s\leq M$.
Since $\frac{\pi}{2}<B<\frac{\pi}{2}$, it follows from $0<\alpha M<\pi$ that
$B+\alpha M=\arctan\left(\cot(\alpha M)-\frac{\lambda\mu^{-1}}{\sin(\alpha M)}\right)+\alpha M<\arctan\left(\cot(\alpha M)\right)+\alpha M,$
because the arctangent is a monotonically increasing function and $\frac{\lambda\mu^{-1}}{\sin(\alpha M)}>0$ when $\alpha M\in(0,\pi)$. Exploiting that $\cot(\alpha M)=\tan(\frac{\pi}{2}-\alpha M)$, we get $\frac{\pi}{2}<B+\alpha M<\frac{\pi}{2}$ for all $0<\alpha M<\pi$. Consequently, $\mu(s)$ is smooth curve that is strictly positive for all $0\leq s\leq M$ and solves the Euler–Lagrange equation. Hence, $\sigma(s)=(\mu(s),\mathbf{v}_{1}(s),\ldots,\mathbf{v}_{d}(s))$ is a geodesic connecting $p$ and $q$. This proves the other direction and concludes the proof.
The foregoing result implies that by changing the $\alpha$-warped metric of $\mathbb{R}_{>0}\times\mathbb{S}^{\mathbf{n},\mathbf{k}}$ we can modify whether two points $p$ and $q$ are connected by a minimizing geodesic. Specifically, we have the following result.
Proposition 4.4.
The pre-Segre–Veronese manifold $\mathcal{P}_{\alpha}^{\mathbf{k}}$ is geodesically connected if and only if $0<\alpha<\frac{1}{\sqrt{k_{1}+\cdots+k_{d}}}$.
Proof 4.5.
Let $p,q\in\mathcal{P}_{\alpha}^{\mathbf{k}}$. Then, $M^{2}(p,q)\leq\sum_{i=1}^{d}k_{i}\pi^{2}$, with equality attained at $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ and $q=(\lambda,-\mathbf{u}_{1},\ldots,-\mathbf{u}_{d})$. The result follows from Proposition 4.2.
4.2 Pre-Segre–Veronese manifolds
Having determined when two points on the pre-Segre–Veronese manifold $\mathcal{P}_{\alpha}^{\mathbf{k}}$ can be connected by a minimizing geodesic, we proceed by computing such a geodesic between two $\alpha$-compatible points.
Theorem 4.6 (Logarithmic map on $\mathcal{P}_{\alpha}^{\mathbf{k}}$).
Let
$p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}\quad\text{and}\quad q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$
be such that $\mathbf{u}_{i}\neq-\mathbf{v}_{i}$ for all $i=1,\dots,d$. Let $M=M(p,q)$ from Definition 4.1. If $p$ and $q$ are $\alpha$-compatible, then
$\operatorname{log}_{p}(q)=\bigl{(}\mu\cos(\alpha M)-\lambda,\dot{\mathbf{u}}_{1},\dots,\dot{\mathbf{u}}_{d}\bigr{)},$
where
$\dot{\mathbf{u}}_{i}:=(\mathbf{v}_{i}-\left\langle{}\mathbf{u}_{i},\mathbf{v}_{i}\right\rangle\mathbf{u}_{i})\frac{\mu\sin(\alpha M)\sphericalangle(\mathbf{u}_{i},\mathbf{v}_{i})}{\lambda\alpha M\sin(\sphericalangle(\mathbf{u}_{i},\mathbf{v}_{i}))},\quad i=1,\dots,d,$
and the expressions $\sin(x)/x$ and $x/\sin(x)$ should be interpreted as $1$ at $x=0$. Especially, if $M=0$, then $\operatorname{log}_{p}(q)=(\mu-\lambda,0,\dots,0).$
Proof 4.7.
We are looking for a geodesic $\gamma$ with starting point $p$ and ending point $q$.
The case $M=0$ follows immediately from Proposition 3.1. Hence, in the remainder of the proof we consider the case $M>0$.
As in the proof of Proposition 3.1, the idea is to use that the spherical part of $\gamma$ is a geodesic on $\mathbb{S}^{n_{1}-1}\times\dots\times\mathbb{S}^{n_{d}-1}$. Since the logarithmic map on a sphere is known, we can pose as ansatz that $\operatorname{log}_{p}(q)$ is of the form
$(\dot{\lambda},\dot{\mathbf{u}}_{1},\dots,\dot{\mathbf{u}}_{d})=(C,D\operatorname{log}_{\mathbf{u}_{1}}(\mathbf{v}_{1}),\dots,D\operatorname{log}_{\mathbf{u}_{d}}(\mathbf{v}_{d}))$
for some $C\in\mathbb{R}$ and $D>0$. Now we just have to solve $q={}\operatorname{exp}_{p}(\operatorname{log}_{p}q)$ for $C$ and $D$.
Recall the expression for $\operatorname{exp}_{p}$ from Proposition 3.1. The first component yields
(18) $(\lambda+C)^{2}+(\lambda\alpha DM)^{2}=\mu^{2},$
having used that $\|\dot{\mathbf{u}}_{i}\|=D\|\log_{\mathbf{u}_{i}}\mathbf{v}_{i}\|=D\sphericalangle(\mathbf{u}_{i},\mathbf{v}_{i})$. Since $M>0$, there exists at least one $1\leq i\leq d$ so that $\mathbf{u}_{i}\neq\mathbf{v}_{i}$. In all of these coordinates, we get
[…]
$\sphericalangle(\mathbf{u}_{i},\mathbf{v}_{i})=a_{i}={}$ $\frac{D\sphericalangle(\mathbf{u}_{i},\mathbf{v}_{i})}{\alpha DM}\left(\frac{\pi}{2}-\arctan\left(\frac{\lambda+C}{\alpha\lambda DM}\right)\right).$
In this way we find
(20) $\mathrm{cotan}(\alpha M)={}$ $\frac{\lambda+C}{\alpha\lambda DM}.$
The square system of polynomial equations Eqs. 18 and 20 in $C$ and $D$ together with the constraint $D>0$ defines the intersection of a circle with a line on the half-plane $D>0$. Ignoring the $D>0$ constraint, the two intersection points are
$C={}\varsigma\mu\cos(\alpha M)-\lambda\quad\text{and}\quad D={}\varsigma\frac{\mu\sin(\alpha M)}{\lambda\alpha M},\quad\text{for }\varsigma\in\{-1,1\}.$
By $\alpha$-compatibility we have $0<\alpha M<\pi$, and since we posited $D>0$, $\varsigma$ has to be 11.
With the above values for $C$ and $D$, one verifies that $\exp_{p}\circ\log_{p}$ is the identity in the first component Eq. 18. The other components of $\exp_{p}\circ\log_{p}$ are also the identity, as can be verified by plugging the spherical logarithmic map [10, Example 10.21]
$\log_{\mathbf{u}_{i}}(\mathbf{v}_{i})={}$ $(\mathbf{v}_{i}-\left\langle{}\mathbf{v}_{i},\mathbf{u}_{i}\right\rangle\mathbf{u}_{i})\frac{\sphericalangle(\mathbf{u}_{i},\mathbf{v}_{i})}{\sin(\sphericalangle(\mathbf{u}_{i},\mathbf{v}_{i}))}$
into Eq. 19. This concludes the proof.
Remark 4.8.
We caution that if $\mathbf{u}_{i}=-\mathbf{v}_{i}$ for some $i$, then the logarithmic map is not well defined. However, $p$ and $q$ can still be connected by a geodesic. It is just that this geodesic is not unique. Choosing $\dot{\mathbf{u}}_{i}$ to be any vector orthogonal to $\mathbf{u}_{i}$ and with norm $\pi$ will produce such a geodesic.
An immediate consequence of the proof of the previous theorem is the following expression for the distance between points on the pre-Segre–Veronese manifold.
Proposition 4.9 (Distance between compatible points on $\mathcal{P}_{\alpha}^{\mathbf{k}}$).
Consider two points $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ that are $\alpha$-compatible. Let $M:=M(p,q)$ from Definition 4.1. Then,
(21) $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)=\sqrt{\lambda^{2}-2\lambda\mu\cos(\alpha M)+\mu^{2}}.$
Proof 4.10.
For $M(p,q)=0$, it follows immediately from Proposition 3.1 that the distance between $p$ and $q$ on $\mathcal{P}_{\alpha}^{\mathbf{k}}$ is equal to $|\absolutevalue{\lambda-\mu}$. This coincides with Eq. 21.
Let $M(p,q)>0$. The proof of Theorem 4.6 shows that $q$ is reached at $s=M(p,q)$ when the geodesic is parameterized to have unit speed on the spheres. Hence, we can use the reparameterization in Eq. 14 to determine the distance $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)$. It follows that it can be expressed as
[…]
where $B=\arctan\left(\frac{\cos(\alpha M)-\lambda\mu^{-1}}{\sin(\alpha M)}\right)$. We proceed by simplifying this formula.
[…]
$\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)$ $=\lambda\cos(B)\frac{\sin(B)\cos(\alpha M)+\cos(B)\sin(\alpha M)}{\cos(B)\cos(\alpha M)-\sin(B)\sin(\alpha M)}-\lambda\sin(B)$
$=\frac{\lambda\sin(\alpha M)}{\cos(B)\cos(\alpha M)-\sin(B)\sin(\alpha M)}.$
[…]
$\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)={}$ $\frac{\lambda\sin(\alpha M)\sqrt{1+\left(\frac{\cos(\alpha M)-\lambda\mu^{-1}}{\sin(\alpha M)}\right)^{2}}}{\cos(\alpha M)-\left(\frac{\cos(\alpha M)-\lambda\mu^{-1}}{\sin(\alpha M)}\right)\sin(\alpha M)}$
[…]
having used that $0<\sin(\alpha M)$ because $0<\alpha M<\pi$. The last equation is equivalent to Eq. 21, concluding the proof.
A formula, equivalent to Eq. 21, is
(22) $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)=\sqrt{(\lambda-\mu)^{2}+4\lambda\mu\sin(\alpha M/2)},$
which is more numerically stable for points that are close together.
The distance between $p$ and $q$ on $\mathcal{P}_{\alpha}^{\mathbf{k}}$ is thus given by the law of cosines applied to an imaginary triangle where one side has length $\lambda$, the other side has length $\mu$, and the angle is the spherical distance $M(p,q)$ multiplied by the warping factor $\alpha$. We thus immediately obtain the following result.
Corollary 4.11.
The distance $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)$ is monotonically increasing as a function of the spherical distance $M(p,q)$ for $0<\alpha M(p,q)<\pi$.
Interestingly, we can also determine the distance between incompatible points, even though minimizing geodesics do not exist between them. While a minimizing geodesic does not exist, there exists a limiting piecewise smooth curve that is the infimizer of the length functional in Eq. 10. In Fig. 1, for example, this limiting piecewise smooth curve is the straight line from $(1,0)$ to $(0,0)$, composed with the straight line from $(0,0)$ to $(0,1)$.
Proposition 4.12 (Distance between incompatible points on $\mathcal{P}_{\alpha}^{\mathbf{k}}$).
Two $\alpha$-incompatible points $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ are at
$\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)=\lambda+\mu.$
Proof 4.13.
Let $\gamma_{n}$ be a sequence of unit speed curves in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ that connect two incompatible points $p$ and $q$, and whose lengths approach the infimum.
To lower bound $\ell(\gamma_{n})$, note that $\gamma_{n}$ must get arbitrarily close to 0 as $n\to\infty$, since otherwise the sequence lives in a compact subset (norm restricted from above and below) of $\mathcal{P}_{\alpha}^{\mathbf{k}}$ and would thus have a (uniformly [^1] Unit speed implies equi-Lipschitz continuous, implies uniform convergence to a Lipschitz continuous curve.) convergent subsequence. The limit of this subsequence would realize the infimum, and thus be a geodesic, contradicting our assumption about $p$ and $q$ being incompatible.
So fix $n$ and let $a=(\epsilon,\mathbf{a}_{1},\dots,\mathbf{a}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ be a point on $\gamma_{n}$ and let $\gamma^{(1)}$ denote the part of $\gamma_{n}$ connecting $p$ and $a$ and let $\gamma^{(2)}$ be the part of $\gamma_{n}$ connecting $a$ and $q$. Then,
$\ell(\gamma_{n})=\ell(\gamma^{(1)})+\ell(\gamma^{(2)}).$
Moreover, if $\gamma^{(1)}(s)=(\nu(s),\mathbf{x}_{1}(s),\dots,\mathbf{x}_{d}(s))$, then
[…]
A similar argument shows that $\ell(\gamma^{(2)})\geq\mu-\epsilon$. Especially, for every $\epsilon$, there exists a $n$ such that $\ell(\gamma_{n})\geq\lambda+\mu-2\epsilon$.
To upper bound $\inf_{\gamma}\ell(\gamma)$, consider the curve that goes first in a straight line from $p=(\lambda,\mathbf{u}_{1},\dots,\mathbf{u}_{d})$ to $(\epsilon,\mathbf{u}_{1},\dots,\mathbf{u}_{d})$, then in a circular arc from $(\epsilon,\mathbf{u}_{1},\dots,\mathbf{u}_{d})$ to $(\epsilon,\mathbf{v}_{1},\dots,\mathbf{v}_{d})$, and then in a straight line from $(\epsilon,\mathbf{u}_{1},\dots,\mathbf{u}_{d})$ to $(\epsilon,\mathbf{v}_{1},\dots,\mathbf{v}_{d})$ to $q=(\mu,\mathbf{u}_{1},\dots,\mathbf{u}_{d})$ to $(\epsilon,\mathbf{v}_{1},\dots,\mathbf{v}_{d})$. It has length $\lambda+\mu+(\alpha M-2)\epsilon$.
Combining these two bounds and letting $\epsilon\to 0$ proves the proposition.
4.3 Segre–Veronese manifolds
By leveraging Lemmas 2.3 and 4.6, minimizing geodesics on the Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ that connect two rank-$1$ tensors are readily obtained. Indeed, a minimizing geodesic connecting $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ is the pushforward of any geodesic whose length coincides with
$\min_{\begin{subarray}{c}p\in\otimes^{-1}(\mathpzc{P}),\,q\in\otimes^{-1}(\mathpzc{Q})\end{subarray}}\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q).$
Corollary 4.11 then leads us naturally to the following terminology.
Definition 4.14 (Matched representatives).
Let $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$. We say that the representatives $p^{\star}\in\otimes^{-1}(\mathpzc{P})$ and $q^{\star}\in\otimes^{-1}(\mathpzc{Q})$ are matched if
$(p^{\star},q^{\star})\in\underset{\begin{subarray}{c}p\in\otimes^{-1}(\mathpzc{P}),\,q\in\otimes^{-1}(\mathpzc{Q})\end{subarray}}{\arg\min}M(p,q),$
where $M$ is the distance from Definition 4.1.
We are now ready to state and prove the main result about minimizing geodesics connecting two rank-$1$ tensors on the $\alpha$-warped Segre–Veronese manifold.
Theorem 4.15 (The logarithmic map of $\mathcal{S}_{\alpha}^{\mathbf{k}}$).
Let $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ be rank-$1$ tensors. If $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\otimes^{-1}(\mathpzc{P})$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\otimes^{-1}(\mathpzc{Q})$ are matched, $\alpha$-compatible representatives, then
$\log{P}\mathpzc{Q}=(\mathrm{d}_{p}\otimes)(\log_{p}q).$
Proof 4.16.
[…]
The following two results follow similarly.
Proposition 4.17.
Two rank-1 tensors $\mathpzc{P}$ and $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ can be connected by a minimizing geodesic in $\mathcal{S}_{\alpha}^{\mathbf{k}}$ if and only if $\mathpzc{P}$ and $\mathpzc{Q}$ have matched $\alpha$-compatible representatives.
Proposition 4.18 (Distance between points on $\mathcal{S}_{\alpha}^{\mathbf{k}}$).
Let $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$. If $p=(\lambda,\mathbf{u}_{1},\dots,\mathbf{u}_{d})\in\otimes^{-1}(\mathpzc{P})$ and $q=(\mu,\mathbf{v}_{1},\dots,\mathbf{v}_{d})\in\otimes^{-1}(\mathpzc{Q})$ are matched, $\alpha$-compatible representatives, then
$\operatorname{dist}_{\mathcal{S}_{\alpha}^{\mathbf{k}}}(\mathpzc{P},\mathpzc{Q})=\operatorname{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q).$
The straightforward approach to determine matched representatives of $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ consists of choosing an arbitrary $p^{\star}\in\otimes^{-1}(\mathpzc{P})$ and then computing all distances $M(p^{\star},q)$ for all $q\in\otimes^{-1}(\mathpzc{Q})$ and selecting a minimum. By Lemma 2.2, the possible $q$ can be labeled by $(\sigma_{1},\dots,\sigma_{d})\in\{\,-1,1\,\}^{d}$ such that $\sigma_{1}^{k_{1}}\cdots\sigma_{d}^{k_{d}}=1$. Thus this strategy has a computational complexity of $\mathcal{O}(2^{d}\dim\mathcal{S}_{\alpha}^{\mathbf{k}})$, where $\mathcal{O}(\dim\mathcal{S}_{\alpha}^{\mathbf{k}})$ is the asymptotic cost for computing one distance $M(p^{\star},q)$. We now explain how a matched representative can be constructed efficiently in $\mathcal{O}(d\dim\mathcal{S}_{\alpha}^{\mathbf{k}})$ operations.
Let $q_{\mathrm{ref}}\in\otimes^{-1}(\mathcal{Q})$ be a reference preimage. We then observe that
$M(p^{\star},q)^{2}=M(p^{\star},q_{\mathrm{ref}})^{2}+\sum_{i=1}^{d}\frac{1-\sigma_{i}}{2}\Delta_{i}.$
with $\Delta_{i}=k_{i}((\sphericalangle_{i}-\pi)^{2}-\sphericalangle_{i}^{2})$, where $\sphericalangle_{i}$ is the angle between the $i$th components of $p^{\star}$ and $q_{\mathrm{ref}}$. Minimizing $M(p^{\star},q)$ thus corresponds to maximizing $\sum_{i=1}^{d}{\sigma_{i}}\Delta_{i}$ over $\sigma_{i}\in\{-1,1\}$. An unconstrained optimum is
(23) $\sigma_{i}=\begin{cases}1,&\text{ if $\Delta_{i}$ is nonnegative},\\ -1,&\text{otherwise}.\end{cases}$
If this assignment also satisfies the constraint $\sigma_{1}^{k_{1}}\cdots\sigma_{d}^{k_{d}}=1$, so that $q\in\otimes^{-1}(q_{\mathrm{ref}})$, then we are done. Otherwise, there is at least one smallest $|\absolutevalue{\Delta_{\ell}}$ among the odd $k_{\ell}$’s. The sign of the corresponding $\sigma_{\ell}$ should be flipped.
Proposition 4.19 (Matchmaking).
The algorithm described in the previous paragraph produces matched representatives.
Proof 4.20.
A detailed proof is given in Section A.3.
The construction of the matchmaking algorithm suggests that determining an analogue of Proposition 4.4 is more complicated for the Segre–Veronese manifold. Essentially, we have to consider the following problem: Given $p^{\star}\in\mathcal{P}_{\alpha}^{\mathbf{k}}$, what is the set of $q\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ such that the corresponding matched representatives $q^{\star}\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ are $\alpha$-compatible with $p^{\star}$, and when is this set the whole of $\mathcal{P}_{\alpha}^{\mathbf{k}}$? By our calculations, this question can be reduced to a combinatorial optimization problem. As we were unable to produce a closed form of its solution, we present only the following general result.
Proposition 4.21.
The Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ is geodesically connected if $0<\alpha<1/\sqrt{k_{1}+\dots+k_{d}}$ and is not geodesically connected if $2/\sqrt{k_{1}+\dots+k_{d}}\leq\alpha$.
Proof 4.22.
The first statement is a corollary of Lemmas 2.2 and 4.4.
To prove the second statement, assume $\alpha\geq 2/\sqrt{k_{1}+\dots+k_{d}}$ and pick $p$, $q\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ with angles $\sphericalangle_{1}=\dots=\sphericalangle_{d}=\frac{\pi}{2}$ between their components. These two representatives are matched. However, using $M(p,q)^{2}=(k_{1}+\dots+k_{d})\frac{\pi^{2}}{4}$ in Definition 4.1 shows that they are not $\alpha$-compatible.
For $1/\sqrt{k_{1}+\dots+k_{d}}<\alpha<2/\sqrt{k_{1}+\dots+k_{d}}$, geodesic connectedness is more complicated and depends also on the parity of the $k_{i}$’s. The full analysis is beyond the scope of this article.
5 Sectional curvature
Formulae for the sectional curvature of a manifold are useful in many situations, for example when approximating or compressing data that lives on said manifold [43, 25, 84], or to ensure uniqueness of the Fréchet mean [46, 47, 3], as required in the experiment from Section 6. We compute these curvatures here for the $\alpha$-warped Segre–Veronese manifolds.
Consider first the $\alpha$-warped Segre manifold $\mathcal{S}_{\alpha}=\mathcal{S}_{\alpha}^{(1,\dots,1)}$. To compute its curvature, we recall the following general result about warped products.
Proposition 5.1 (O’Neill [69, 7.42]).
For $(a,b)\in\mathcal{M}\times_{f}\mathcal{N}$, let $\dot{\mathbf{x}}$, $\dot{\mathbf{y}}$, $\dot{\mathbf{z}}\in\mathrm{T}_{a}\mathcal{M}$ and $\dot{\mathbf{u}}$, $\dot{\mathbf{v}}$, $\dot{\mathbf{w}}\in\mathrm{T}_{b}\mathcal{N}$ be tangent vectors, which naturally lift to $\mathrm{T}_{(a,b)}(\mathcal{M}\times_{f}\mathcal{N})$. Let $R_{\mathcal{M}}$ and $R_{\mathcal{N}}$ be the Riemann tensors on $\mathcal{M}$ and $\mathcal{N}$, respectively. Then, the Riemann tensor $R$ on $\mathcal{M}\times_{f}\mathcal{N}$ is described by: [^2] Note that we use a definition of $R$ that is consistent with [53] rather than [69]. The result is that some signs differ in our expressions from theirs.
  1. 1.
    $R(\dot{\mathbf{x}},\dot{\mathbf{y}})\dot{\mathbf{z}}=R_{\mathcal{M}}(\dot{\mathbf{x}},\dot{\mathbf{y}})\dot{\mathbf{z}}$,
  2. 2.
    $R(\dot{\mathbf{u}},\dot{\mathbf{x}})\dot{\mathbf{y}}=-\frac{1}{f}\operatorname{Hess}[f](\dot{\mathbf{x}},\dot{\mathbf{y}})\dot{\mathbf{u}}$,
  3. 3.
    $R(\dot{\mathbf{x}},\dot{\mathbf{y}})\dot{\mathbf{u}}=0$,
  4. 4.
    $R(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{x}}=0$,
  5. 5.
    $R(\dot{\mathbf{x}},\dot{\mathbf{u}})\dot{\mathbf{v}}=-\frac{\left\langle{}\dot{\mathbf{u}},\dot{\mathbf{v}}\right\rangle}{f}\nabla_{\dot{\mathbf{x}}}(\operatorname{grad}f)$,
  6. 6.
    $R(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{w}}=R_{\mathcal{N}}(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{w}}+\frac{\norm{\operatorname{grad}f}^{2}}{f^{2}}\left(\left\langle{}\dot{\mathbf{u}},\dot{\mathbf{w}}\right\rangle\dot{\mathbf{v}}-\left\langle{}\dot{\mathbf{v}},\dot{\mathbf{w}}\right\rangle\dot{\mathbf{u}}\right)$.
Herein, $\left\langle{}\cdot,\cdot\right\rangle$ is the inner product on $\mathrm{T}_{(a,b)}(\mathcal{M}\times_{f}\mathcal{N})$, $\operatorname{grad}f$ is the Riemannian gradient of $f$, $\operatorname{Hess}[f]$ is the Riemannian Hessian of $f$, and $\nabla$ is the Levi–Civita connection.
For $\mathcal{S}_{\alpha}$, which locally looks like the warped product manifold $\mathcal{P}_{\alpha}$, these expressions simplify considerably. Applying Proposition 5.1 to $\mathcal{N}=\mathbb{S}^{{\mathbf{n}},\mathbf{1}}=\mathbb{S}^{{n_{1}}-1}\times\dots\times\mathbb{S}^{{n_{d}}-1}$ and $\mathcal{M}=\mathbb{R}_{>0}$ while observing that $\operatorname{grad}f=\alpha$ and $R_{\mathcal{M}}=0$ yields the next result.
Corollary 5.2.
In the notation of Proposition 5.1, the Riemann tensor $R$ of $\mathcal{S}_{\alpha}$ at $\lambda\mathbf{u}_{1}\otimes\dots\otimes\mathbf{u}_{d}$ satisfies
$R(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{w}}=R_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{w}}+\frac{1}{\lambda^{2}}(\left\langle{}\dot{\mathbf{u}},\dot{\mathbf{w}}\right\rangle\dot{\mathbf{v}}-\left\langle{}\dot{\mathbf{v}},\dot{\mathbf{w}}\right\rangle\dot{\mathbf{u}}),$
and it is 0 in the remaining directions.
As sectional curvatures can be defined in terms of the Riemann tensor, we quickly find the next consequence of the last result.
Corollary 5.3.
For $(\mathbf{u}_{1},\dots,\mathbf{u}_{d})\in\mathbb{S}^{{\mathbf{n}},\mathbf{1}}$, let the lifts of the tangent vectors $\dot{\mathbf{u}}_{i}\in\mathrm{T}_{\mathbf{u}_{i}}\mathbb{S}^{{n_{i}}-1}$ and $\dot{\mathbf{v}}_{j}\in\mathrm{T}_{\mathbf{u}_{j}}\mathbb{S}^{{n_{j}}-1}$ be orthonormal with respect to $\langle\cdot,\cdot\rangle_{\mathcal{P}_{\alpha}}$. The sectional curvature $K$ of $\mathcal{S}_{\alpha}$ at $\lambda\mathbf{u}_{1}\otimes\dots\otimes\mathbf{u}_{d}$ satisfies
$K(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})=\begin{cases}\frac{1-\alpha^{2}}{\alpha^{2}\lambda^{2}},&\text{if $i=j$},\\ -\frac{1}{\lambda^{2}},&\text{otherwise},\end{cases}$
and $K=0$ in all remaining directions.
Proof 5.4.
Recall that $\left\langle{}\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j}\right\rangle_{\mathcal{P}_{\alpha}}=\alpha^{2}\lambda^{2}\left\langle{}\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j}\right\rangle_{\mathbb{S}^{{\mathbf{n}},\mathbf{1}}}$. We compute
$K(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})={}$ $\left\langle{}R(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})\dot{\mathbf{v}}_{j},\dot{\mathbf{u}}_{i}\right\rangle_{\mathcal{P}_{\alpha}}$
$={}$ $\left\langle{}R_{\mathbb{S}^{{\mathbf{n}},\mathbf{1}}}(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})\dot{\mathbf{v}}_{j},\dot{\mathbf{u}}_{i}\right\rangle_{\mathcal{P}_{\alpha}}+\frac{1}{\lambda^{2}}\left(\left\langle{}\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j}\right\rangle_{\mathcal{P}_{\alpha}}^{2}-\left\langle{}\dot{\mathbf{v}}_{j},\dot{\mathbf{v}}_{j}\right\rangle_{\mathcal{P}_{\alpha}}\left\langle{}\dot{\mathbf{u}}_{i},\dot{\mathbf{u}}_{i}\right\rangle_{\mathcal{P}_{\alpha}}\right)$
$={}$ $\alpha^{2}\lambda^{2}\left\langle{}R_{\mathbb{S}^{{\mathbf{n}},\mathbf{1}}}(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})\dot{\mathbf{v}}_{j},\dot{\mathbf{u}}_{i}\right\rangle_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}-\frac{1}{\lambda^{2}}$
$={}$ $\frac{\alpha^{2}\lambda^{2}}{\alpha^{4}\lambda^{4}}\left\langle{}R_{\mathbb{S}^{{\mathbf{n}},\mathbf{1}}}(\alpha\lambda\dot{\mathbf{u}}_{i},\alpha\lambda\dot{\mathbf{v}}_{j})\alpha\lambda\dot{\mathbf{v}}_{j},\alpha\lambda\dot{\mathbf{u}}_{i}\right\rangle_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}-\frac{1}{\lambda^{2}}$
$={}$ $\frac{1}{\alpha^{2}\lambda^{2}}K_{\mathbb{S}^{{\mathbf{n}},\mathbf{1}}}(\alpha\lambda\dot{\mathbf{u}}_{i},\alpha\lambda\dot{\mathbf{v}}_{j})-\frac{1}{\lambda^{2}}$
$={}$ $\frac{1}{\alpha^{2}\lambda^{2}}\delta_{ij}-\frac{1}{\lambda^{2}}.$
In all other directions the sectional curvature is 0, because $R=0$ in these directions.
The Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$, $\mathbf{k}=(k_{1},\dots,k_{d})$, is a Riemannian submanifold of the Segre manifold $\mathcal{S}_{\alpha}^{\mathbf{1}}$, where $\mathbf{1}$ is a vector of $k_{1}+\dots+k_{d}$ ones. If we look closely at the expression in Proposition 3.1 for geodesics on $\mathcal{S}_{\alpha}^{\mathbf{1}}$, we see that starting out in a tangent direction to $\mathcal{S}_{\alpha}^{\mathbf{k}}\subset\mathcal{S}_{\alpha}^{\mathbf{1}}$ produces a path wholly contained in $\mathcal{S}_{\alpha}^{\mathbf{k}}$. We have thus proved the following result.
Proposition 5.5.
The Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ is a totally geodesic submanifold of the Segre manifold $\mathcal{S}_{\alpha}^{\mathbf{1}}$, and so its sectional curvature is described by restricting $K$ in Corollary 5.3 to the appropriate tangent subspace.
6 Numerical experiment
The $\alpha$-warped Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ was implemented in Julia, including connecting and starting geodesics and the geodesic distance. Our implementation is available as part of the Manifolds.jl library [5].
As a numerical example illustration of the connecting geodesics in the $\alpha$-warped geometry of the Segre manifold, we recall the consensus aggregation problem from Chafamo, Shanmugam, and Tokcan [16]: Given $M$ approximate tensor rank decompositions of the same tensor $\mathpzc{A}\in\mathbb{R}^{n_{1}\times\dots\times n_{d}}$,
$\mathpzc{A}\approx\mathpzc{A}^{(m)}:=\sum_{i=1}^{r}\mathpzc{A}_{i}^{(m)}:=\sum_{i=1}^{r}\mathbf{a}_{1i}^{(m)}\otimes\dots\otimes\mathbf{a}_{di}^{(m)},\quad m=1,\,\dots,\,$M,
estimate the true decomposition of $\mathpzc{A}$.
To solve the above problem, a consensus aggregation algorithm was proposed in [16] that takes elementwise medians of corresponding vectors $\mathbf{a}_{{l}i}^{(m)}$, $m=1$, …, $M$, of matching summands in the decomposition. While tensor rank decompositions usually have generically unique decompositions into a set of rank-$1$ tensors [61, 81, 6], to perform aggregation one needs to determine how these rank-$1$ tensors match up across multiple decompositions. For this reason, the rank-$1$ tensors are first matched using a ($k$-means) clustering algorithm in [16]. This strategy can work well if the decomposition problem is well-conditioned [12], so that $\mathpzc{A}^{(m)}$ being close to $\mathpzc{A}$ will imply that the corresponding rank-$1$ tensors are close too. After matching up the rank-$1$ tensors, the results are aggregated in [16] by setting
$\widehat{\mathbf{a}}_{i}^{l}={}$ $\operatorname{median}(\mathbf{a}^{(m)}_{{l}i}:m=1,\,\dots,\,M),\quad i=1,\ldots,r,\;{l}=1,\ldots,d.$
The median is taken elementwise. Note that it is possible for $\widehat{\mathbf{a}}_{i}^{l}$ to have different values if different representatives are chosen. This is circumvented in [16] by aggregating positive data, for which there is a natural, positive, representative. The aggregated tensor decomposition is then
$\mathpzc{A}\approx\widehat{\mathpzc{A}}:=\sum_{i=1}^{r}\widehat{\mathbf{a}}_{i}^{1}\otimes\dots\otimes\widehat{\mathbf{a}}_{i}^{d}.$
The above is a simplified description of the essential ingredients of [16, Section 3].
From a geometric viewpoint, it is natural to estimate each rank-$1$ tensor with the Fréchet mean on the warped Segre manifold $\mathcal{S}_{\alpha}$. Recall that the Fréchet mean or Riemannian center of mass of points $x_{1},\dots,x_{k}$ on a Riemannian manifold $\mathcal{M}$ is
$\operatorname{mean}_{\mathcal{M}}(x_{1},\dots,x_{k})=\underset{p\in\mathcal{M}}{\arg\min}\sum_{i=1}^{n}\mathrm{dist}_{\mathcal{M}}(p,p_{i})^{2}.$
Its uniqueness has been investigated among others in [46, 47, 3]. For manifolds with negative sectional curvatures, the Fréchet mean is unique and the above defines a valid function. Hence, we propose to estimate $\mathpzc{A}$ instead as
[…]
To guarantee that all points are comparable so that the Fréchet mean is well defined, we choose $\alpha=1/\sqrt{k_{1}+\dots+k_{d}}-\sqrt{\epsilon}$ where $\epsilon$ is the machine precision of double-precision floating-point numbers. Proposition 4.21 then guarantees all points are comparable.
We reproduce the experiment from [16, Section 4.1], using the authors’ code with the same parameter values and initialization seed. Figure 2 compares our estimate to their estimate and to the baseline of doing no consensus aggregation. The baseline, ZIPTF, consists of just doing one zero-inflated Poisson tensor factorization as described in [16, Section 3.2]. C-ZIPTF is the consensus aggregation of $20$ ZIPTF’s by Chafamo, Shanmugam, and Tokcan’s proposed method of medians. F-ZIPTF is similarly the consensus aggregation of $20$ ZIPTF’s by our proposed method of Fréchet means. The Fréchet mean is approximated by successive geodesic interpolation, see for example [19, 17, 18, 22]. We conclude that the accuracy of both consensus variants is better than the baseline. There does not appear to be a meaningful difference between C-ZIPTF and F-ZIPTF. This nevertheless validates the correctness and utility of connecting geodesics for the $\alpha$-warped Segre manifold.
Figure 2: We recreate the setup from [16, Section 4.1] and compare the Riemannian distance between the approximate decompositions and the ground truth.
The Julia source code for the experiments is available publicly [42].
Appendix A Proof of the technical lemmas
A.1 Proof of Lemma 2.2
Each $\imath_{\sigma}$ is an isometry because $\mathcal{P}_{\alpha}^{\mathbf{k}}$ is a product manifold where $\imath_{\sigma}$ acts only on the spheres and the maps $\mathrm{Id}$ and $\mathrm{Id}$ are basic isometries of the sphere with the standard Euclidean inner product, and hence also with any scaled inner product.
Normality, i.e., $\otimes(p)=\otimes(q)$ implies there exists a $\sigma$ such that $q=\imath_{\sigma}(p)$ [69, Appendix A], follows from the multilinearity of the tensor product $\otimes$ [33, Chapter 1].
By the same arguments as [7, Section 4.1], $\otimes$ is a smooth covering map. Hence $\dim\mathcal{P}_{\alpha}^{\mathbf{k}}=\dim\mathcal{S}_{\alpha}^{\mathbf{k}}$ and the differential $\mathrm{d}\otimes:\mathrm{T}\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathrm{T}\mathcal{S}_{\alpha}^{\mathbf{k}}$ is everywhere left-invertible.
The key part is showing that $\otimes:\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathcal{S}_{\alpha}^{\mathbf{k}}$ is Riemannian, i.e., the metric of $\mathcal{S}_{\alpha}^{\mathbf{k}}$ is the pushforward of the one of $\mathcal{P}_{\alpha}^{\mathbf{k}}$, which is a straightforward computation. Let $\dot{\mathbf{u}}_{i}\in\mathrm{T}_{\mathbf{u}_{i}}\mathbb{S}^{n_{1}-1}=\mathbf{u}_{i}^{\perp}$ and $\dot{\lambda}\in\mathbb{R}$ be arbitrary. By differentiating $\otimes:\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathcal{S}_{\alpha}^{\mathbf{k}}$, we get
$(\mathrm{d}_{(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})}\otimes)(\dot{\lambda},\dot{\mathbf{u}}_{1},\ldots,\dot{\mathbf{u}}_{d})\\ =\dot{\lambda}\mathpzc{U}+\lambda(\dot{\nu}_{k_{1}}(\dot{\mathbf{u}}_{1})\otimes\mathbf{u}_{2}^{\otimes k_{2}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}})+\cdots+\lambda(\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d-1}^{\otimes k_{d-1}}\otimes\dot{\nu}_{k_{d}}(\dot{\mathbf{u}}_{d})),$
where $\mathpzc{U}$ and $\nu_{k}(\mathbf{u})$ are as in Eqs. 16 and 17. The $\alpha$-warped metric on $\mathcal{S}_{\alpha}^{\mathbf{k}}$ satisfies
$\varsigma^{\alpha}_{\lambda\mathpzc{U}}\left(\mathrm{d}\otimes(\dot{x},\dot{\mathbf{u}}_{1},\ldots,\dot{\mathbf{u}}_{d}),\mathrm{d}\otimes(\dot{y},\dot{\mathbf{v}}_{1},\ldots,\dot{\mathbf{v}}_{d})\right)=\dot{x}\dot{y}+(\alpha\lambda)^{2}\langle\dot{\mathpzc{U}},\dot{\mathpzc{V}}\rangle,$
having dropped the subscript $(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ of the differential and where
$\dot{\mathpzc{U}}$ $=(\dot{\nu}_{k_{1}}(\dot{\mathbf{u}}_{1})\otimes\mathbf{u}_{2}^{\otimes k_{2}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}})+\cdots+(\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d-1}^{\otimes k_{d-1}}\otimes\dot{\nu}_{k_{d}}(\dot{\mathbf{u}}_{d})),$
$\dot{\mathpzc{V}}$ $=(\dot{\nu}_{k_{1}}(\dot{\mathbf{v}}_{1})\otimes\mathbf{u}_{2}^{\otimes k_{2}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}})+\cdots+(\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d-1}^{\otimes k_{d-1}}\otimes\dot{\nu}_{k_{d}}(\dot{\mathbf{v}}_{d})).$
Recall from [36] that the Euclidean inner product between rank-$1$ tensors satisfies $\langle\mathbf{u}_{1}\otimes\cdots\otimes\mathbf{u}_{d},\mathbf{v}_{1}\otimes\cdots\otimes\mathbf{v}_{d}\rangle=\prod_{i=1}^{d}\langle\mathbf{u}_{i},\mathbf{v}_{i}\rangle.$ Hence, rank-$1$ tensors are orthogonal to one another if they are orthogonal in at least one factor. We compute that $\langle\nu_{k_{i}}(\dot{\mathbf{u}}_{i}),\mathbf{u}_{i}^{\otimes k}\rangle=0,$ while $\langle\nu_{k_{i}}(\dot{\mathbf{u}}_{i}),\nu_{k_{i}}(\dot{\mathbf{v}}_{i})\rangle=k_{i}\langle\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{i}\rangle$ if $\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{i}\in\mathrm{T}_{\mathbf{u}_{i}}\mathbb{S}^{n}$. As a consequence, we can observe that the cross terms in $\langle\dot{\mathpzc{U}},\dot{\mathpzc{V}}\rangle$ vanish and we are left with
$\langle\dot{\mathpzc{U}},\dot{\mathpzc{V}}\rangle=\sum_{i=1}^{d}\langle\dot{\nu}_{k_{i}}(\dot{\mathbf{u}}_{i}),\dot{\nu}_{k_{i}}(\dot{\mathbf{v}}_{i})\rangle=\sum_{i=1}^{d}k_{i}\langle\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{i}\rangle.$
This shows that $\varsigma_{\lambda\mathpzc{U}}$ is the pushforward of $g_{\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d}}$ under $\otimes$, concluding the proof.
A.2 Proof of Lemma 2.3
Every smooth curve on $\mathcal{N}$ has such a unique smooth lift through a normal covering by [69, Lemma A.9]. Moreover, as $\phi$ is a local isometry that restricts to an isometry on evenly covered neighborhoods, the length of $\gamma$ and its unique lift $\widetilde{\gamma}$ in $\mathcal{M}$ coincide. Indeed, if we partition the lift $\widetilde{\gamma}$ into the curve segments $\widetilde{\gamma}_{i}$ such that $\widetilde{\gamma}_{i}\subset\mathcal{M}_{i}\subset\mathcal{M}$ with $\phi|_{\mathcal{M}_{i}}:\mathcal{M}_{i}\to\mathcal{N}_{i}$ an isometry (such a partition exists and is used to prove the existence of a unique lift in [69, Lemma A.9]), then
$\ell(\widetilde{\gamma})=\sum_{i=1}^{k}\ell(\widetilde{\gamma}_{i})=\sum_{i=1}^{k}\ell(\phi(\widetilde{\gamma}_{i}))=\ell(\gamma).$
This concludes the first part of the proof.
If there were a strictly shorter smooth curve $\widetilde{\gamma}^{\prime}\subset\mathcal{M}$ connecting the endpoints of the lift $\widetilde{\gamma}$, then by the same partitioning argument $\phi\circ\widetilde{\gamma}^{\prime}\subset\mathcal{N}$ would be a shorter piecewise smooth curve ($\phi$ is a local embedding) than $\gamma$, contradicting its minimality. This proves the second part.
A.3 Proof of Proposition 4.19
We want to maximize $f=\sum_{i=1}^{d}\sigma_{i}\Delta_{i}$ over $\sigma_{i}\in\set{-1,1}$ with the constraint that $\sigma_{1}^{k_{1}}\dots\sigma_{d}^{k_{d}}=1$.
We can assume without loss of generality that all $k_{i}$’s are odd. Indeed, even $k_{i}$ do not impact the constraint, so we can choose the corresponding sign so that $|\Delta_{i}|=\sigma_{i}\Delta_{i}$, which is clearly optimal, also without constraint, because of the bilinear structure of $f$. The matching algorithm chooses these signs precisely in this way.
We can further assume without loss of generality that all $\Delta_{i}$ are nonzero, for otherwise we could take the unconstrained optimum Eq. 23 and simply swap the sign of a $\sigma_{j}$ for which $\Delta_{j}=0$ to obtain a feasible solution that is also globally optimal, which is what the matching algorithm does.
Let $(\varsigma_{1},\dots,\varsigma_{d})$ be a constrained optimizer of $f$ that differs in the least number of indices from the unconstrained optimizer $(\sigma_{1},\dots,\sigma_{d})$ from Eq. 23. Denote the number of different indices by $p=\sharp\{i\in\{1,\dots,d\}\mid\sigma_{i}\neq\varsigma_{i}\}$. Then, there are three cases:
  1. $p=0$.
    The unconstrained optimum given by Eq. 23 and chosen by the algorithm satisfies the constraint.
  2. $p=1$.
    If $f_{0}$ denotes the value of $f$ for the unconstrained optimum, such a constrained optimum has value $|f_{0}-2\absolutevalue{\Delta_{i}}$. Hence, $i$ must correspond to the minimal $|\absolutevalue{\Delta_{i}}$, which the matching algorithm indeed selects by construction.
  3. $p\geq 2$.
    Let $i\neq j$ be any two indices for which $\varsigma_{i}\neq\sigma_{i}$ and $\varsigma_{j}\neq\sigma_{j}$. By changing the signs of $\varsigma_{i}$ and $\varsigma_{j}$ to match those of $\sigma_{i}$ and $\sigma_{j}$, respectively, this increases the objective by $2|\Delta_{i}|+2|\Delta_{j}|$. Since $|\Delta_{i}|,|\Delta_{j}|>0$, this modified, feasible constrained point with p2p-2 different indices has a strictly higher objective value than $(\varsigma_{1},\dots,\varsigma_{d})$, which was supposed to be the constrained minimizer with minimal $p$. This is a contradiction, so this case cannot occur.
This concludes the proof.
A.4 Proof of Eq. 14
Let $C=\frac{\dot{\lambda}}{\alpha\lambda M}$, and recall that $B=\mathrm{arctan}(C)$ and $A=\lambda\cos(B)=\frac{\lambda}{\sqrt{1+C^{2}}}.$ Define
(24) $z(t):=\frac{\lambda+\dot{\lambda}t}{\alpha\lambda Mt}=\frac{1}{\alpha Mt}+C,\quad\text{so that}\quad z^{\prime}(t)=-\frac{1}{\alpha Mt^{2}}.$
In this notation, Eq. 14 becomes $\alpha s(t)=\frac{\pi}{2}-\arctan(z(t)),$ so that on the one hand
$\alpha s^{\prime}(t)=-\frac{z^{\prime}(t)}{1+z^{2}(t)}.$
Since $\sqrt{\dot{\lambda}^{2}+(\alpha\lambda M)^{2}}=\alpha\lambda M\sqrt{1+C^{2}}$, Eq. 13 simplifies as follows:
$\alpha s^{\prime}(t)$ $=\cos^{2}(\alpha s(t)+B)A^{-1}\sqrt{\dot{\lambda}^{2}+(\alpha\lambda M)^{2}}$
$=\cos^{2}(\alpha s(t)+B)\alpha M(1+C^{2})$
$=\alpha M(1+C^{2})\cos^{2}\Bigl{(}\frac{\pi}{2}-(\arctan(z(t))-\arctan(C))\Bigr{)}$
$=\alpha M(1+C^{2})\sin^{2}(\arctan(z(t))-\arctan(C)).$
Using basic trigonometric identities, it is easy to obtain the next identity:
(25) $\sin(\arctan(x)-\arctan(y))=\frac{x-y}{\sqrt{1+x^{2}}\sqrt{1+y^{2}}}.$
This allows us to conclude the proof because
$\alpha s^{\prime}(t)=\alpha M(1+C^{2})\frac{(z(t)-C)^{2}}{(1+z^{2}(t))(1+C^{2})}=\alpha M\frac{(\alpha Mt)^{-2}}{1+z^{2}(t)}=-\frac{z^{\prime}(t)}{1+z^{2}(t)}.$
A.5 Proof of Eq. 15
Using the same notation as in Section A.4, we get that $\alpha s(1)=\frac{\pi}{2}-\arctan(z(1)),$ so that
$\gamma(1)=\mu(s(1))=\frac{\lambda\cos(B)}{\cos(\alpha s(1)+B)}$ $=\frac{\lambda}{\sqrt{1+C^{2}}}\cdot\frac{1}{\sin(\arctan(z(1))-\arctan(C))}$
$=\frac{\lambda}{\sqrt{1+C^{2}}}\frac{\sqrt{1+C^{2}}\sqrt{1+z^{2}(1)}}{z(1)-C}$
$=\alpha\lambda M\sqrt{1+z^{2}(1)},$
where we used Eq. 25 in the penultimate and Eq. 24 in the final equality. Since $z^{2}(1)=(\lambda+\dot{\lambda})^{2}(\alpha\lambda M)^{-2}$, the proof is concluded.