### v3

Abstract
Segre–Veronese manifolds are smooth submanifolds of tensors comprising the partially symmetric rank-$1$ tensors. We investigate a one-parameter family of warped geometries of Segre–Veronese manifolds, which includes the standard Euclidean geometry. This parameter controls by how much spherical tangent directions are weighted relative to radial tangent directions. We present closed expressions for the exponential map, the logarithmic map, and the intrinsic distance in these warped Segre–Veronese manifolds, which can be computed efficiently numerically. It is shown that Segre–Veronese manifolds are not geodesically connected in the Euclidean geometry, while they are for some values of the warping parameter. The benefits of geodesic connectedness may outweigh using the Euclidean geometry in certain applications. One such application is presented: numerically computing the Riemannian center of mass for averaging rank-$1$ tensors.
keywords:
Segre–Veronese manifold; partially symmetric rank-$1$ tensors; geodesic; exponential map; logarithmic map; warped geometry
{AMS}
15A69; 65F99; 53C22; 53A35; 14N07
1 Introduction
The $\mathbf{k}$th Segre–Veronese manifold $\mathcal{S}^{\mathbf{k}}$ of partially symmetric rank-$1$ tensors in the space of $n_{1}^{k_{1}}\times n_{2}^{k_{2}}\times\cdots\times n_{d}^{k_{d}}$ real arrays is a fundamental set of elementary tensors [51]. Special cases include the punctured plane
$\mathcal{S}^{(1)}:=\mathbb{R}^{n}_{*}:=\set{\lambda\mathbf{u}\mid\lambda\in\mathbb{R}_{*},\norm{\mathbf{u}}=1}$;
the manifold of symmetric rank-$1$ matrices
$\mathcal{S}^{(2)}:=\set{\lambda\mathbf{u}\mathbf{u}^{\mathsf{T}}\mid\lambda\in\mathbb{R}_{*},\,\norm{\mathbf{u}}=1}\subset\mathbb{R}^{n\times n},$
and, more generally, the Veronese manifold $\mathcal{S}^{(d)}$ of symmetric rank-$1$ tensors [51, 23]; and the manifold of rank-$1$ matrices
$\mathcal{S}^{(1,1)}:=\set{\lambda\mathbf{u}\mathbf{v}^{\mathsf{T}}\mid\lambda\in\mathbb{R}_{*},\,\norm{\mathbf{u}}=1,\,\norm{\mathbf{v}}=1}\subset\mathbb{R}^{m\times n},$
and, more generally, the Segre manifold $\mathcal{S}^{(1,\ldots,1)}$ of rank-$1$ tensors [51]. Analogously to the role of rank-$1$ matrices in a low-rank matrix decomposition, Segre–Veronese manifolds are the basic building blocks of more advanced tensor rank decomposition models [51]. Such sums of partially symmetric rank-$1$ tensors feature in a plethora of applications in biomedical engineering, chemometrics, computer science, data science, machine learning, psychometrics, signal processing, and statistics, among others; see the overview articles [50, 75, 69, 4, 62, 9] and the references therein.
The algebraic geometry of the Segre–Veronese manifold and its secant varieties have been thoroughly studied since the $19$th century [8]. By contrast, comparatively little is known about the Riemannian geometry of Segre–Veronese manifolds, let alone of (the smooth loci of) its higher secant varieties. To our knowledge, only the Euclidean geometry of the Segre manifold of rank-$1$ tensors (i.e., $\mathbf{k}=(1,\dots,1)$) as a submanifold of $\mathbb{R}^{n_{1}\times\dots\times n_{d}}$ was investigated in the prior works [78, 77].
The present paper studies essential properties of the Segre–Veronese manifold $\mathcal{S}^{\mathbf{k}}$ as a metric space. We are primarily interested in characterizing its geodesics. These curves $\gamma\subset\mathcal{S}^{\mathbf{k}}$ can be thought of as (locally) length-minimizing curves that generalize straight lines in Euclidean spaces. Being solutions of differential equations, geodesics can be characterized as the solution of an initial value or a boundary value problem:
  1. 1.
[…]
  2. 2.
    Two “endpoints” $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}^{\mathbf{k}}$ can be specified to identify a geodesic $\gamma$ that satisfies $\gamma(0)=\mathpzc{P}$ and $\gamma(\ell)=\mathpzc{Q}$, where $\ell$ is the distance between $\mathpzc{P}$ and $\mathpzc{Q}$. The map that takes these two endpoints as input and sends them to the corresponding geodesic is called the logarithmic map of the manifold.
We will characterize the geodesics of Segre–Veronese manifolds by presenting explicit expressions for their exponential and logarithmic maps.
1.1 Applications of geodesics
The logarithmic map finds application in (i) statistics on manifolds [30], (ii) interpolation on manifolds [84], and (iii) approximation of maps into manifolds. First, the definition and computation of statistics on manifolds often involves geodesics connecting two points [30]. The simplest statistic, the Fréchet mean [31], Riemannian center of mass, or Karcher mean [45], in a manifold $\mathcal{M}$ consists of the point(s) that minimize the integral of the squared intrinsic distances to a fixed set of points in $\mathcal{M}$. The gradient of the squared intrinsic distance function from $x$ to a fixed point $x^{\star}\in\mathcal{M}$ is the tangent vector at $x$ of a geodesic connecting $x$ and $x^{\star}$. Therefore, geodesics appear for computing such a Fréchet mean both in gradient-based Riemannian optimization algorithms [71, 70, 44] and other approximation algorithms [45, 44, 61, 19]. The logarithmic map also appears in the definition and computation of higher-order statistics on manifolds, such as principal geodesic analysis [27, 28, 76] and the Gaussian distribution and its parameter estimation [70, 18]. It is an essential ingredient in geodesic regression [11, 29, 30] and generalizations thereof [40]. Second, many algorithms for interpolation of data in manifolds rely on the logarithmic map [67, 59, 43, 72, 66, 32, 73, 83, 84, 65, 86, 85]. Third, geodesics connecting endpoints are an essential building block of algorithms for constructing an approximation of a function from a Euclidean space into a Riemannian manifold. They are used in the Riemannian moving least squares algorithm [34, 35, 74], and the pullback-based methods of [56, 42, 82].
The exponential map has applications in (i) Riemannian optimization [1, 10], and (ii) integration on manifolds [37]. First, the exponential map (or approximations thereof called retractions [79]) are an essential component of Riemannian optimization algorithms, which optimize a smooth objective function over a constraint set that is a smooth manifold [1, 10]. Second, the exponential map, the projection retraction [2], or other retractions are essential in intrinsic integration schemes for vector fields on manifolds, such as the ones arising from differential equations on manifolds [24, 39, 64, 63, 38, 54, 37] or the projected ones from dynamic low-rank approximation methods (see [48, 49, 58, 81, 57, 26, 15, 41, 20]).
The foregoing discussion described general potential applications of geodesics. Next, we highlight plausible applications in the context of Segre–Veronese manifolds. Geodesics can be used in Riemannian optimization for seeking an approximate partially symmetric tensor rank decomposition [8]. Such algorithms were already applied to approximate tensor rank decompositions using a retraction in [13] and using the exponential map in [78], and to symmetric tensor rank or Waring decompositions [8, 23] using a retraction in [47]. As illustrated in Section 6, computing a Fréchet mean is a natural alternative to averaging tensor rank decompositions based on factor matrices as was proposed in [16] to improve the robustness of stochastic tensor decompositions.
1.2 Contributions
Our main contributions are analytic expressions of the exponential and logarithmic maps of Segre–Veronese manifolds, in Corollary 3.4 and Theorem 4.15 respectively. We establish a few auxiliary results, such as a formula for the distance between points in Segre–Veronese manifolds in Proposition 4.18.
Proposition 4.17 entails that not every pair of partially symmetric rank-$1$ tensors in a Segre–Veronese manifold can be connected by a minimizing geodesic in the standard Euclidean geometry. In some applications, such as the one in Section 6, it is more important that every pair of points is connected by a minimizing geodesic than that Euclidean geometry is used. For this reason, we investigate $\alpha$-warped geometries [21] of Segre–Veronese manifolds. These geometries are described in detail in Section 2; essentially, they scale, by a factor of $\alpha>0$, the length of all tangent directions that do not infinitesimally change the Euclidean norm. The choice $\alpha=1$ corresponds to the Euclidean geometry. Figure 1 illustrates several geodesic segments, specified either by an initial point and a tangent vector as in the left panel, or by two “endpoints” as in the right panel; that is, these can be interpreted as the geodesics captured by, respectively, the exponential and logarithmic maps.
(a) Geodesics starting at $p=(0,1)$ in the tangent direction $\dot{p}=(1,-1)$ in several $\alpha$-warped geometries. All geodesic segments have the same length $\sqrt{2}$ in their $\alpha$-warped geometry, while their Euclidean lengths are seen to be very different. In the $\alpha$-warped geometries with $\alpha<1$, the geodesics will first spiral towards the circle about the origin of radius $\alpha$ and then spiral out again from it. This will be evident from Propositions 3.1 and 3.4.
(b) Minimizing geodesic segments between $p=(0,1)$ and $q=(1,0)$ in several $\alpha$-warped geometries. As $\alpha\to 0$, the geodesic segment will approach the unit circle arc that connects $p$ and $q$. As $\alpha\to\infty$, the geodesic segment will converge to the piecewise linear path formed by the line segments from $(0,1)$ to $(1,0)$ via the node $(0,0)$. This will be evident from Theorems 4.6, 4.15 and 4.12. We need a bit more text to get to the next line.
Figure 1: Geodesics in the $\alpha$-warped geometry of the punctured plane $\mathcal{S}^{(1)}$ for $\alpha=0.05$ (the red paths curving towards the origin, i.e., acceleration vectors pointing toward the origin), $\frac{1}{4}$, $\frac{1}{2}$, $\frac{3}{4}$, 11 (the purple straight lines between $(0,1)$ and $(1,0)$), $\frac{5}{4}$, $\frac{3}{2}$, $\frac{7}{4}$, and $1.95$ (the blue paths curving away from the origin, i.e., acceleration vectors pointing away from $(0,0)$). Essentially, the left panel shows evaluations of the exponential map, and the right panel shows evaluations of the logarithmic map.
1.3 Outline
The next section introduces background material from Riemannian geometry and describes $\alpha$-warped geometries of $\mathbb{R}_{*}^{n}$. General $\alpha$-warped Segre–Veronese manifolds are obtained as Riemannian embedded submanifolds thereof. We also give an alternative characterization of $\alpha$-warped Segre–Veronese manifolds, revealing them to be covered by a warped product of the positive real numbers and a product of spheres, which we call a pre-Segre–Veronese manifold. This characterization is exploited in the subsequent sections: the exponential maps of (pre-)Segre–Veronese manifolds are computed in Section 3 and their logarithmic maps are computed in Section 4. We present the sectional curvatures of Segre–Veronese manifolds in Section 5. Thereafter, in Section 6, we detail an example application of geodesics for computing the Fréchet mean.
Notation
For convenience, scalars will be typeset in a lower case, vectors in a bold lower case, tensors in a calligraphic upper case, and manifolds in another calligraphic upper case. We will emphasize that a vector lives in a tangent space by adding an upper dot, e.g., $\dot{\mathbf{v}}$.
Unless otherwise specified, $(\mathbb{R}^{n},\langle\cdot,\cdot\rangle)$ will be abbreviated to $\mathbb{R}^{n}$; it is the inner product space of $n$-dimensional real vectors equipped with the standard Euclidean inner product. If $\mathcal{M}\subset\mathbb{R}^{n}$ is any subset, then $\mathbb{S}(\mathcal{M}):=\set{\mathbf{u}\in\mathcal{M}\mid\norm{\mathbf{u}}=1}$ is the space of vectors of $\mathcal{M}$ of unit Euclidean norm. We use the shorthand $\mathbb{S}^{n-1}:=\mathbb{S}(\mathbb{R}^{n})$ for the unit sphere in $\mathbb{R}^{n}$. The angular distance from $\mathbf{u}$ to $\mathbf{v}$, i.e., the arclength of a shortest great circle segment connecting them, is $\sphericalangle(\mathbf{u},\mathbf{v})=\cos^{-1}(\langle\mathbf{v},\mathbf{u}\rangle)\in[0,\pi]$.
Let $\mathcal{M}$ be a manifold. Its tangent space at $x\in\mathcal{M}$ is denoted by $\mathrm{T}_{x}\mathcal{M}$. The exponential map of $\mathcal{M}$ at $x$ is $\exp_{x}\colon\Omega_{x}\to\mathcal{M}$, where $\Omega_{x}$ is a suitable open neighborhood of $0\in\mathrm{T}_{x}\mathcal{M}$. The logarithmic map of $\mathcal{M}$ at $x$ is $\log_{x}\colon\mathcal{M}_{x}\to\mathrm{T}_{x}\mathcal{M}$, where $\mathcal{M}_{x}$ is a suitable open neighborhood of $x\in\mathcal{M}$. For a differentiable curve $\gamma\colon\mathbb{R}\to\mathcal{M}$, we denote the derivative to its parameter (see [53, Section 3] for a definition) using a prime, i.e.,
$\gamma^{\prime}(t):=\frac{\mathrm{d}}{\mathrm{d}t}\gamma(t)\quad\in\mathrm{T}_{x(t)}\mathcal{M}.$
The exponential and logarithmic maps of a sphere $\mathbb{S}^{n-1}$ are, respectively,
[…]
2 $\alpha$-warped geometries for Segre–Veronese manifolds
The next subsection recalls preliminary results from Riemannian geometry. In particular, the warped product of manifolds will be described. Then, in Section 2.2, the $\mathbf{k}$th Segre–Veronese manifold is formally defined and presented as a Riemannian embedded submanifold of a warped geometry of the space of tensors. Section 2.3 shows how Segre–Veronese manifolds can be realized as a normal Riemannian covering of a simpler warped product involving only spheres and the positive reals.
2.1 Warped geometries of $\mathbb{R}^{n}_{*}$
Recall that a smooth embedded manifold $\mathcal{M}\subset\mathbb{R}^{N}$ equipped with a Riemannian metric—an inner product $g_{x}\colon\mathrm{T}_{x}\mathcal{M}\times\mathrm{T}_{x}\mathcal{M}\to\mathbb{R}$ on its tangent spaces $\mathrm{T}_{x}\mathcal{M}$ that satisfies some technical regularity conditions, see, e.g., [52]—induces an intrinsic metric structure on $\mathcal{M}$ as follows. A Riemannian metric induces a norm $\norm{\dot{\mathbf{x}}}_{x}:=\sqrt{g_{x}(\dot{\mathbf{x}},\dot{\mathbf{x}})}$ for all $\dot{\mathbf{x}}\in\mathrm{T}_{x}\mathcal{M}$, which can be used to measure the length $\ell$ of piecewise smooth curves $\gamma\colon[0,1]\to\mathcal{M}$:
[…]
Letting the set of all piecewise smooth curves be our length structure $L$, the manifold $\mathcal{M}$ becomes a length space [14]. The intrinsic metric is then
$\mathrm{dist}_{\mathcal{M}}(x,y):=\inf_{\begin{subarray}{c}\gamma\in L,\\ \gamma(0)=x,\,\gamma(1)=y\end{subarray}}\ell(\gamma).$
A diffeomorphism (a smooth bijection with smooth inverse) between manifolds that preserves the intrinsic metric is called an isometry.
A piecewise smooth curve $\gamma\in L$ connecting $x=\gamma(0)$ and $y=\gamma(1)$ such that the length of $\gamma$ equals the distance $\mathrm{dist}_{\mathcal{M}}(x,y)$ is called a minimizing geodesic [52]. That is, a minimizing geodesic is a shortest path in $L$ between $x$ and $y$ contained in $\mathcal{M}$. More generally, a geodesic is any piecewise smooth curve that satisfies the first-order optimality condition, i.e., is a critical point, of the variational problem $\min_{\gamma\in L}\ell(\gamma)$ subject to appropriate constraints (such as fixing two endpoints). Such curves are always locally length-minimizing [52, Theorem 6.12] in the sense that they are minimizing geodesics on a segment $\gamma|_{[0,t)}$ for some $t>0$. Note that isometries map geodesics to geodesics.
For each tangent vector $\dot{\mathbf{v}}\in\mathrm{T}_{x}\mathcal{M}$, there is a unique constant-speed geodesic satisfying $\gamma(0)=x$, $\gamma^{\prime}(0)=\dot{\mathbf{v}}$ [52, Theorem 4.10]. We can build a map from (an open subset of) $\mathrm{T}_{x}\mathcal{M}$ to $\mathcal{M}$ via $\dot{\mathbf{v}}\mapsto\gamma(1)$; this is the exponential map. It is a diffeomorphism between a neighborhood of $0\in\mathrm{T}_{x}\mathcal{M}$ and a neighborhood of $x\in\mathcal{M}$ [52, Lemma 5.10]. Its inverse is called the logarithmic map.
We also mention two important concepts in Riemannian geometry: The Riemann curvature tensor and the sectional curvature. The Riemann curvature tensor on a Riemannian manifold $\mathcal{M}$ is a multilinear map $R\colon\mathrm{T}_{p}\mathcal{M}\times\mathrm{T}_{p}\mathcal{M}\times\mathrm{T}_{p}\mathcal{M}\to\mathrm{T}_{p}\mathcal{M}$ that measures the change in a tangent vector as it is parallel transported around a small loop. The sectional curvature is a map $K\colon\mathrm{T}_{p}\mathcal{M}\times\mathrm{T}_{p}\mathcal{M}\to\mathbb{R}$ that measures the angular defect of small triangles around $p$. See [52, Chapter 7] for precise definitions.
The punctured space $\mathbb{R}^{n}_{*}$ is an open submanifold of $\mathbb{R}^{n}$; hence, it is a smooth manifold. It can be equipped with the standard Euclidean metric $g$ at $\mathbf{a}\in\mathbb{R}^{n}_{*}$, so that $g_{\mathbf{a}}(\dot{\mathbf{x}},\dot{\mathbf{y}})=\langle\dot{\mathbf{x}},\dot{\mathbf{y}}\rangle=\dot{\mathbf{x}}^{\mathsf{T}}\dot{\mathbf{y}},$ where $\dot{\mathbf{x}}$, $\dot{\mathbf{y}}\in\mathrm{T}_{\mathbf{a}}\mathbb{R}^{n}_{*}=\mathbb{R}^{n}$ and $\langle\cdot,\cdot\rangle$ is the standard Euclidean inner product. This leads to the Euclidean geometry of $(\mathbb{R}^{n}_{*},g)$. Other meaningful geometries of $\mathbb{R}^{n}_{*}$ exist. For example, $\mathbb{R}^{n}_{*}$ is diffeomorphic to the product of the positive real line and an (n1)(n-1)-dimensional sphere: $\mathbb{R}^{n}_{*}\simeq\mathbb{R}_{>0}\times\mathbb{S}^{n-1}.$ Under this identification, the above standard Euclidean metric can be expressed as
$g_{\lambda\mathbf{u}}(\dot{x}\mathbf{u}+\lambda\dot{\mathbf{v}},\dot{y}\mathbf{u}+\lambda\dot{\mathbf{w}})=\dot{x}\dot{y}+\lambda^{2}\langle\dot{\mathbf{v}},\dot{\mathbf{w}}\rangle,$
where $\dot{x}$, $\dot{y}\in\mathbb{R}=\mathrm{T}_{\lambda}\mathbb{R}_{>0}$ and $\dot{\mathbf{v}}$, $\dot{\mathbf{w}}\in\mathbf{u}^{\perp}=\mathrm{T}_{\mathbf{u}}\mathbb{S}^{n-1}$. Herein, the inner products on the right are the inner products inherited from the ambient Euclidean spaces $\mathbb{R}$ and $\mathbb{R}^{n}$, respectively. Note in particular that the metric that $\mathbb{R}^{n}_{*}$ inherits from $\mathbb{R}^{n}$ is not equal to the product metric, $\dot{x}\dot{y}+\left\langle{}\dot{\mathbf{v}},\dot{\mathbf{w}}\right\rangle$, of $\mathbb{R}_{>0}$ and $\mathbb{S}^{n-1}$. Instead, it is a so-called warped product [68, Chapter 7] with the identity map as warping function. Recall from [68, Chapter 7] that if we are given Riemannian manifolds $(\mathcal{M},g)$ and $(\mathcal{N},h)$, then we can construct their warped product with warping function $f\colon\mathcal{M}\to\mathbb{R}$ as $\mathcal{M}\times_{f}\mathcal{N}:=(\mathcal{M}\times\mathcal{N},g\times_{f}h),$ where
$(g\times_{f}h)_{(m,n)}((\dot{\mathbf{m}}_{1},\dot{\mathbf{n}}_{1}),(\dot{\mathbf{m}}_{2},\dot{\mathbf{n}}_{2}))=g_{m}(\dot{\mathbf{m}}_{1},\dot{\mathbf{m}}_{2})+f^{2}(m)\cdot h_{n}(\dot{\mathbf{n}}_{1},\dot{\mathbf{n}}_{2}).$
Rather than considering only the Euclidean geometry of $\mathbb{R}^{n}_{*}$, we also consider its $\alpha$-warped geometries in this paper; that is,
(3) $\mathcal{R}_{\alpha}=(\mathbb{R}^{n}_{*},\varsigma^{\alpha}):=\mathbb{R}_{>0}\times_{\alpha\mathrm{Id}}\mathbb{S}^{n-1}\quad\text{with }\alpha>0,$
whose warped product metric satisfies
(4) $\varsigma^{\alpha}_{\lambda\mathbf{u}}(\dot{x}\mathbf{u}+\lambda\dot{\mathbf{v}},\dot{y}\mathbf{u}+\lambda\dot{\mathbf{w}})=\dot{x}\dot{y}+(\alpha\lambda)^{2}\langle\dot{\mathbf{v}},\dot{\mathbf{w}}\rangle.$
If $\alpha=1$, then the $1$-warped geometry is the Euclidean geometry of $\mathbb{R}^{n}_{*}$, with straight lines as geodesics between points. If $\alpha>1$, then distances in the spherical directions will be elongated, implying that geodesics between points will tend to be forced toward the origin. On the other hand, if $0<\alpha<1$, then spherical movements are easier, implying that geodesics between points will bend away from the origin. This is visualized in Fig. 1(b).
2.2 The $\alpha$-warped Segre–Veronese submanifold
The punctured space $\mathbb{R}^{n}_{*}$ is the most basic example (with $\mathbf{k}=(1)$) of the $\mathbf{k}$th Segre–Veronese manifold
(5) $\mathcal{S}^{\mathbf{k}}_{\pm}:=\set{\lambda\mathbf{x}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{x}_{d}^{\otimes k_{d}}\mid\lambda\in\mathbb{R}_{*},\;\mathbf{x}_{i}\in\mathbb{S}^{n_{i}-1},\,i=1,\ldots,d}$,
where $\mathbf{x}_{i}^{\otimes k_{i}}$ is the $k_{i}$-fold tensor product of $\mathbf{x}_{i}$ with itself, i.e., $\mathbf{x}_{i}^{\otimes k_{i}}:=\mathbf{x}_{i}\otimes\cdots\otimes\mathbf{x}_{i}$. Recall that the tensor product of vectors in $\mathbb{R}^{n_{j}}$, $j=1$, …, $d$, can be embedded naturally into the space of $n_{1}\times\cdots\times n_{d}$ arrays, as follows:
$(\mathbf{x}_{1}\otimes\cdots\otimes\mathbf{x}_{d})(i_{1},\ldots,i_{d}):=\mathbf{x}_{1}(i_{1})\cdots\mathbf{x}_{d}(i_{d}),\qquad 1\leq i_{j}\leq n_{j},\,j=1,\ldots,d,$
where $\mathbf{x}_{j}\in\mathbb{R}^{n_{j}}$ and $\mathbf{x}_{j}(i_{j})$ denotes the $i_{j}$th coefficient of $\mathbf{x}_{j}$ with respect to a basis of $\mathbb{R}^{n_{j}}$; see, e.g., [55]. That is, the coordinates of a tensor product are obtained by multiplying the coordinates of all vectors in all possible ways.
Let us also define the ?positive? Segre–Veronese manifold
(6) $\mathcal{S}^{\mathbf{k}}:=\set{\lambda\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}}\mid\lambda\in\mathbb{R}_{>0},\;\mathbf{u}_{i}\in\mathbb{S}^{n_{i}-1},\,i=1,\ldots,d}$,
which is a smooth embedded submanifold (see e.g. [51, Section 4.3] and [53, Chapter 5]) that is diffeomorphic to $\mathbb{R}_{>0}\times\mathbb{S}(\mathcal{S}^{\mathbf{k}}).$ We can thus equip it with the $\alpha$-warped metric that it inherits as a submanifold of $\mathcal{R}_{\alpha}=\mathbb{R}_{>0}\times\mathbb{S}^{N-1}$ from Eq. 3, where $N=n_{1}^{k_{1}}\cdots n_{d}^{k_{d}}$. The resulting Riemannian submanifold will be denoted by
(7) $\mathcal{S}_{\alpha}^{\mathbf{k}}:=\mathbb{R}_{>0}\times_{\alpha\mathrm{Id}}\mathbb{S}(\mathcal{S}^{\mathbf{k}})=(\mathcal{S}^{\mathbf{k}},\varsigma^{\alpha})\subset\mathcal{R}_{\alpha},$
and we will refer to it as the $\alpha$-warped Segre–Veronese manifold.
Remark 2.1.
If all $k_{i}$’s are even, then $\mathcal{S}_{\pm}^{\mathbf{k}}$ consists of two connected components, namely the positive Segre–Veronese manifold $\mathcal{S}^{\mathbf{k}}$ and its mirror image along the origin, namely the ?negative? Segre–Veronese manifold $\mathcal{S}^{\mathbf{k}}$. This means that $p\in\mathcal{S}^{\mathbf{k}}$ and $q\in-\mathcal{S}^{\mathbf{k}}$ are not connected by any curve in $\mathcal{S}^{\mathbf{k}}_{\pm}$. In particular, there will be no geodesics connecting them. Nevertheless, the two connected components are isometric, as $\varsigma_{(-\lambda)\mathbf{u}}^{\alpha}=\varsigma_{\lambda\mathbf{u}}^{\alpha}$ (herein, the metric should be interpreted as being defined on $\mathbb{R}_{\neq 0}$), and we have $\mathcal{S}^{\mathbf{k}}=\mathbb{R}_{<0}\times_{\alpha\mathrm{Id}}\mathbb{S}^{N-1}$. As these two connected components are isometric manifolds, their geometry is the same.
On the other hand, if at least one of the $k_{i}$’s is odd, then it is easy to show that $\mathcal{S}_{\pm}^{\mathbf{k}}=\mathcal{S}^{\mathbf{k}}=-\mathcal{S}^{k}$. That is, the ?regular?, ?positive?, and ?negative? flavors of the Segre–Veronese manifold all coincide.
Therefore, we will henceforth study the component $\mathcal{S}^{\mathbf{k}}$ without loss of generality.
2.3 The $\alpha$-warped pre-Segre–Veronese manifold
We study the geometry of $\mathcal{S}_{\alpha}^{\mathbf{k}}$ through the lens of the tensor product $\otimes$. By definition, $\mathcal{S}^{\mathbf{k}}$ is the image of
$\otimes\colon\mathbb{R}_{>0}\times\mathbb{S}^{n_{1}-1}\times\cdots\times\mathbb{S}^{n_{d}-1}$ $\longrightarrow\mathbb{R}^{n_{1}^{k_{1}}\times\cdots\times n_{d}^{k_{d}}}_{*},$
$(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ $\longmapsto\lambda\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}}.$
As the domain is a product manifold, its tangent space at $(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ is
$\mathrm{T}_{\lambda}\mathbb{R}_{>0}\times\mathrm{T}_{\mathbf{u}_{1}}\mathbb{S}^{n_{1}-1}\times\cdots\times\mathrm{T}_{\mathbf{u}_{d}}\mathbb{S}^{n_{d}-1}=\mathbb{R}\times\mathbf{u}_{1}^{\perp}\times\cdots\times\mathbf{u}_{d}^{\perp}.$
We can equip it with the $\alpha$-warped product metric
(8) $g_{(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})}\bigl((\dot{x},\dot{\mathbf{u}}_{1},\ldots,\dot{\mathbf{u}}_{d}),(\dot{y},\dot{\mathbf{v}}_{1},\ldots,\dot{\mathbf{v}}_{d})\bigr):=\dot{x}\dot{y}+(\alpha\lambda)^{2}\sum_{i=1}^{d}k_{i}\langle\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{i}\rangle.$
The resulting Riemannian manifold will be called the $\alpha$-warped pre-Segre–Veronese manifold, and it will be denoted by
(9) $\mathcal{P}_{\alpha}^{\mathbf{k}}=(\mathbb{R}_{>0},\langle\cdot,\cdot\rangle)\times_{\alpha\mathrm{Id}}\underbrace{((\mathbb{S}^{n_{1}-1},k_{1}\langle\cdot,\cdot\rangle)\times\cdots\times(\mathbb{S}^{n_{d}-1},k_{d}\langle\cdot,\cdot\rangle))}_{\mathbb{S}^{\mathbf{n},\mathbf{k}}}.$
Note that the standard Euclidean inner product on the $i$th sphere is weighted by a factor $k_{i}$. The geometry of such a sphere $(\mathbb{S}^{n_{i}-1},k_{i}\langle\cdot,\cdot\rangle)$ is equivalent to the usual Euclidean geometry of the sphere $(\mathbb{S}^{n_{i}-1}_{\sqrt{k_{i}}},\langle\cdot,\cdot\rangle)$ of radius $\sqrt{k_{i}}$.
The key result is that the tensor product takes the pre-Segre–Veronese manifold $\mathcal{P}_{\alpha}^{\mathbf{k}}$ in a standard way to the Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$. Specifically, it is a normal Riemannian covering map [68], which means that (i) every point $\mathpzc{T}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ has the same number, $P$, of preimages in $\mathcal{P}_{\alpha}^{\mathbf{k}}$, (ii) every point $\mathpzc{T}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ has a neighborhood $U_{\mathpzc{T}}$ such that its preimage $\otimes^{-1}(U_{\mathpzc{T}})$ is a disjoint union of $P$ domains, and (iii) on each of those domains, $\otimes$ is an isometry. Property (iii) is called local isometry.
A diffeomorphism $\iota\colon\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathcal{P}_{\alpha}^{\mathbf{k}}$ that satisfies $\otimes\circ\iota=\otimes$ is called a deck transformation. Intuitively, it shuffles the preimages of the covering map. A deck transform that is also an isometry is called an isometric deck transform.
Lemma 2.2.
The map
$\otimes\colon\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathcal{S}_{\alpha}^{\mathbf{k}},\quad(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\mapsto\lambda\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}}$
is a normal Riemannian covering where the isometric deck transforms are of the form $\imath_{\sigma}(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})=(\lambda,\sigma_{1}\mathbf{u}_{1},\ldots,\sigma_{d}\mathbf{u}_{d})$ with $\sigma_{i}\in\set{-1,1}$ and $\sigma_{1}^{k_{1}}\cdots\sigma_{d}^{k_{d}}=1$.
An elementary but crucial fact about normal Riemannian coverings concerns the lengths of curves in the covered manifold. The next result enables the computation of geodesics in $\mathcal{S}_{\alpha}^{\mathbf{k}}$ by studying the geodesics of the warped product $\mathcal{P}_{\alpha}^{\mathbf{k}}$.
Lemma 2.3.
[…]
In particular, a minimizing geodesic segment in $\mathcal{N}$ lifts to a minimizing geodesic segment in $\mathcal{M}$.
These lemmas are proved in Appendix A.
By Lemma 2.3, if we want to find a minimizing geodesic that connects $\mathpzc{P}$ and $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$, it suffices to enumerate the geodesics connecting their preimages in the $\alpha$-warped pre-Segre–Veronese manifold and pick the shortest one. This is essentially what our proposed matchmaking algorithm will do in Proposition 4.19.
The geodesics of $\alpha$-warped product manifolds in general have been studied by Chen [21]. In particular, [21, Lemma 4.3] states that geodesics in the warped product $\mathcal{M}\times_{f}\mathcal{N}$ have the property that their projection onto the second factor $\mathcal{N}$ also yields a geodesic, albeit not parameterized to be of constant speed. We can effectively leverage this result to parameterize geodesic segments passing through a given point in the pre-Segre–Veronese manifold.
Lemma 2.4.
Let $p=(\lambda,\mathbf{u}_{1},\dots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ be a point in the $\alpha$-warped pre-Segre–Veronese manifold. If $\sigma\subset\mathcal{P}_{\alpha}^{\mathbf{k}}$ is a geodesic passing through $p$, then there exist $\eta\in\mathbb{R}_{>0}$, $B\in\mathbb{R}$, and tangent vectors $\dot{\mathbf{v}}_{i}\in\mathrm{T}_{\mathbf{u}_{i}}\mathbb{S}^{n_{i}-1}$ for $i=1,\dots,d$ such that
(10a) $\sigma(s)=\left(\mu(s),\mathbf{v}_{1}(s),\dots,\mathbf{v}_{d}(s)\right),\quad s\in[0,\eta),$
parameterizes a geodesic segment with $\sigma(0)=p$ and either
$\sigma(s)=(\lambda+s,\mathbf{u}_{1},\dots,\mathbf{u}_{d})$
if all $\dot{\mathbf{v}}_{i}=0$, or
  1. 1.
    $\sigma(s)$ has unit spherical speed, i.e.,
    (10b) $\norm{(\mathbf{v}_{1}^{\prime}(s),\dots,\mathbf{v}_{d}^{\prime}(s))}_{(\mathbf{v}_{1}(s),\ldots,\mathbf{v}_{d}(s))}=1;$
  2. 2.
    the real curve is
    (10c) $\mu(s)=\lambda\frac{\cos(B)}{\cos(\alpha s+B)};$
  3. 3.
    the spherical curves are great circles, i.e.,
    (10d) $\mathbf{v}_{i}(s)=\exp^{\mathbb{S}}_{\mathbf{u}_{i}}(s\dot{\mathbf{v}}_{i}),\quad i=1,\ldots,d,$
    where $\exp^{\mathbb{S}}_{\mathbf{u}_{i}}$ is the exponential map of the sphere $\mathbb{S}^{n_{i}-1}$ as in Eq. 1.
Proof 2.5.
Since $\mathcal{P}_{\alpha}^{\mathbf{k}}=\mathbb{R}_{>0}\times_{\alpha\mathrm{Id}}\mathbb{S}^{\mathbf{n},\mathbf{k}}$ is a warped product, Chen’s result [21, Lemma 4.3] implies that a geodesic in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ projects to a geodesic of $\mathbb{S}^{\mathbf{n},\mathbf{k}}$. The geometry of $\mathbb{S}^{\mathbf{n},\mathbf{k}}$ is the standard product geometry of the Euclidean spheres $\mathbb{S}^{n_{i}-1}_{\sqrt{k_{i}}}$ of radius $\sqrt{k_{i}}$. Hence, the geodesics of $\mathbb{S}^{\mathbf{n},\mathbf{k}}$ are the products of the geodesics in the individual spheres $\mathbb{S}^{n_{i}-1}_{\sqrt{k_{i}}}$, which are all great circles. This shows that there exist tangent vectors $\dot{\mathbf{v}}_{i}$ so that all $\mathbf{v}_{i}(s)$’s are of the form Eq. 10d.
The coefficient $\mu\colon\mathbb{R}\to\mathbb{R}$ in Eq. 10a is determined by solving the Euler–Lagrange equations of the length functional, which in the $\alpha$-warped geometry becomes
[…]
In the remainder of the proof, we treat the case where some $\dot{\mathbf{v}}_{i}\neq 0$. Then, we can assume $\sigma$ has been parameterized so that Eq. 10b holds. The critical points of the Euler–Lagrange equation are the solutions of $\frac{\partial L}{\partial\mu}=\frac{\partial}{\partial s}\frac{\partial L}{\partial\mu^{\prime}}$, where $L(s,\mu,\mu^{\prime})=\sqrt{\mu^{\prime}(s)^{2}+\alpha^{2}\mu(s)^{2}}$. This ultimately leads to the differential equation $\alpha^{2}\mu(s)^{2}+2\mu^{\prime}(s)^{2}-\mu(s)\mu^{\prime\prime}(s)=0$ whose solution is of the form $\mu(s)=\frac{A}{\cos(\alpha s+B)}.$ Solving for the constant $A$ by exploiting the starting conditions $\sigma(0)=p$, so $\mu(0)=\lambda$, we arrive at $A=\lambda\cos(B)$. This proves the form Eq. 10 holds.
Note that the characterization of geodesics in Lemma 2.4 is not complete in the sense there are unspecified quantities ($\eta,B,\dot{\mathbf{v}}_{1},\dots,\dot{\mathbf{v}}_{d}$). This is as expected, because geodesics are solutions to the Euler–Lagrange differential equation. Unique solutions to this second-order differential equation require two interpolation conditions, such as passing through a point and its tangent vector (yielding the exponential map), or passing through two points (yielding the logarithmic map).
3 The exponential map
Before presenting the main original contribution in Section 4, we briefly investigate the exponential maps of (pre-)Segre–Veronese manifolds. The following result is a generalization of [78, Theorem 1] and its proof, which dealt only with the usual Euclidean geometry (i.e., $\alpha=1$) of the Segre manifold (i.e., $\mathbf{k}=(1,\ldots,1)$).
Proposition 3.1 (Exponential map of $\mathcal{P}_{\alpha}^{\mathbf{k}}$).
Let
$p=(\lambda,\mathbf{u}_{1},\dots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}\quad\text{and}\quad\dot{p}=(\dot{\lambda},\dot{\mathbf{u}}_{1},\dots,\dot{\mathbf{u}}_{d})\in\mathrm{T}_{p}\mathcal{P}_{\alpha}^{\mathbf{k}}.$
Then, the exponential map satisfies
[…]
where
$N=\sqrt{\sum_{i=1}^{d}k_{i}\norm{\dot{\mathbf{u}}_{i}}^{2}}\quad\text{and}\quad a=\frac{1}{\alpha N}\left(\frac{\pi}{2}-\arctan\left(\frac{\lambda+\dot{\lambda}}{\lambda\alpha N}\right)\right).$
If $N=0$ and $\lambda<\dot{\lambda}$, the geodesics are straight lines:
$\exp_{p}(\dot{p})=(\lambda+\dot{\lambda},\mathbf{u}_{1},\ldots,\mathbf{u}_{d}).$
Proof 3.2.
Recall that the exponential map evaluated at $\dot{p}$ is equivalent to evaluating the constant-speed geodesic $\gamma(t)=(\lambda(t),\mathbf{u}_{1}(t),\dots,\mathbf{u}_{d}(t))$ in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ with starting conditions $\gamma(0)=p$ and $\gamma^{\prime}(0)=\dot{p}$ at $t=1$.
If $N=0$, the form of the geodesic, and hence exponential map, follows immediately from Lemma 2.4. So we only need to consider the case $N>0$.
[…]
$\dot{\mathbf{u}}_{i}=\mathbf{v}_{i}^{\prime}(0)=s^{\prime}(0)\dot{\mathbf{v}}_{i},$
where $\dot{\mathbf{v}}_{i}$ are the tangent vectors from Lemma 2.4. Moreover, $\norm{\mathbf{v}_{i}^{\prime}(0)}_{\mathbf{v}_{i}(0)}s^{\prime}(0)=\sqrt{k_{i}}\norm{\dot{\mathbf{u}}_{i}}$, $i=1,\dots,d$, from which we find that $s^{\prime}(0)=N$, where $N$ is as in the statement of the proposition. Solving for the constant $B$ from Lemma 2.4 by exploiting the starting condition $\mu^{\prime}(0)=\dot{\lambda}$, we arrive at $B=\arctan(\dot{\lambda}/(\alpha\lambda N))$.
Let $A=\lambda\cos(B)$. We have that
[…]
Since $\sigma(s)$ is injective, the parameterization $s(t)$ is a monotone increasing function. Hence, $s^{\prime}(t)\geq 0$. By plugging $\norm{\gamma^{\prime}(t)}_{\gamma(t)}=\norm{\dot{p}}_{p}=\sqrt{\dot{\lambda}^{2}+(\alpha\lambda N)^{2}}$ into Eq. 12, we obtain the differential equation
(13) $s^{\prime}(t)=\cos^{2}(\alpha s(t)+B)\frac{\sqrt{\dot{\lambda}^{2}+(\alpha\lambda N)^{2}}}{\alpha A}.$
Combining this with the initial condition $s(0)=0$, we have the solution
(14) $s(t)=\alpha^{-1}\left(\frac{\pi}{2}-\arctan\left(\frac{\lambda+\dot{\lambda}t}{\alpha\lambda Nt}\right)\right).$
Verifying that this solves Eq. 13 is a straightforward computation; see Section A.4.
To obtain the exponential map, we now just have to evaluate $\gamma(1)$. For the first component of $\gamma(1)$ we find, after some computations (see Section A.5), that
(15) $\mu(s(1))={}$ $\sqrt{(\lambda+\dot{\lambda})^{2}+(\alpha\lambda N)^{2}}.$
The other components $\mathbf{v}_{i}(s(1))$ are obtained immediately by plugging the expression of $s(t)$ from Eq. 14 into Eq. 10d. This concludes the proof.
From this result, it immediately follows that the exponential map is well defined almost everywhere.
Corollary 3.3.
The domain of the exponential map $\exp_{p}$ of $\mathcal{P}_{\alpha}^{\mathbf{k}}$ is the whole tangent space $\mathrm{T}_{p}\mathcal{P}_{\alpha}^{\mathbf{k}}$, except for the half-line $\set{(\dot{\lambda},0,\ldots,0)\mid\dot{\lambda}\leq-\lambda}$.
Having computed the geodesics in the pre-Segre–Veronese manifold, an application of Lemmas 2.2 and 2.3 enables us to push them forward with $\otimes$ to the Segre–Veronese manifold. Note that because Riemannian coverings are local isometries, we are only guaranteed that geodesics are mapped to geodesics, but not that minimizing geodesics are mapped to minimizing geodesics, i.e., the converse of Lemma 2.3 is not true. In this way, we arrive at the geodesics and the following characterization of the exponential map of $\mathcal{S}_{\alpha}^{\mathbf{k}}$.
Corollary 3.4 (The exponential map of $\mathcal{S}_{\alpha}^{\mathbf{k}}$).
Let $\mathpzc{P}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ be a rank-$1$ tensor and let $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\otimes^{-1}(\mathpzc{P})$ be any of its representatives in $\mathcal{P}_{\alpha}^{\mathbf{k}}$. Let $\dot{\mathpzc{P}}\in\mathrm{T}{P}\mathcal{S}_{\alpha}^{\mathbf{k}}$, which can be expressed uniquely as
$\dot{\mathpzc{P}}=\dot{\lambda}\mathpzc{U}+\lambda\left(\nu_{k_{1}}(\dot{\mathbf{u}}_{1})\otimes\mathbf{u}_{2}^{\otimes k_{2}}\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}}+\cdots+\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d-1}^{\otimes k_{d-1}}\otimes\nu_{k_{d}}(\dot{\mathbf{u}}_{d})\right),$
where
(16) $\mathpzc{U}$ $=\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}},$
(17) $\nu_{k_{i}}(\dot{\mathbf{u}}_{i})$ $=(\dot{\mathbf{u}}_{i}\otimes\mathbf{u}_{i}^{\otimes(k_{i}-1)})+\cdots+(\mathbf{u}_{i}^{\otimes(k_{i}-1)}\otimes\dot{\mathbf{u}}_{i})\quad\text{for all }\dot{\mathbf{u}}_{i}\in\mathrm{T}_{\mathbf{u}_{i}}\mathbb{S}^{n}.$
Let $\dot{p}=(\dot{\lambda},\dot{\mathbf{u}}_{1},\ldots,\dot{\mathbf{u}}_{d})\in\mathrm{T}_{p}\mathcal{P}_{\alpha}^{\mathbf{k}}$ be the corresponding lifted tangent vector. Then,
$\exp{P}(\dot{\mathpzc{P}})=\otimes(\exp_{p}(\dot{p})),$
insofar as $\dot{p}$ is in the domain of $\exp_{p}$.
Another interesting consequence of Propositions 3.1 and 3.3 is that it entails the next result about the radius of the largest open ball in $\mathrm{T}{P}\mathcal{S}_{\alpha}^{\mathbf{k}}$ for which $\exp$ is a diffeomorphism, i.e., the local injectivity radius of the Segre–Veronese manifold.
Corollary 3.5 (Injectivity radius).
The local injectivity radius of $\mathcal{P}_{\alpha}^{\mathbf{k}}$, and thus also of $\mathcal{S}_{\alpha}^{\mathbf{k}}$, at $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ is $\lambda$. The (global) injectivity radius of $\mathcal{P}_{\alpha}^{\mathbf{k}}$ and $\mathcal{S}_{\alpha}^{\mathbf{k}}$, i.e., the infimum of the local injectivity radii over the manifold, is zero.
Note that the latter claim is also immediate from the fact that neither $\mathcal{P}_{\alpha}^{\mathbf{k}}$ nor $\mathcal{S}_{\alpha}^{\mathbf{k}}$ are complete metric spaces. Indeed, the sequence $p_{n}=(n^{-1},\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ tends to $(0,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\not\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ as $n\to\infty$, and the image of $p_{n}$ in $\mathcal{S}_{\alpha}^{\mathbf{k}}$ tends to $0\notin\mathcal{S}_{\alpha}^{\mathbf{k}}$.
Finally, we highlight that the exponential map of the pre-Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ can be computed numerically very efficiently by straightforward implementation of the formulas in Proposition 3.1. The cost is about $\mathcal{O}(\sum_{i=1}^{d}n_{i})$ operations to compute $N$ and all $\mathbf{u}_{i}(t)$’s, and a constant number of operations for $\lambda(t)$ and $g(t)$.
4 The logarithmic map
We now compute the logarithmic map of the $\alpha$-warped (pre-)Segre–Veronese manifold, significantly extending prior results of one of the authors, Lars Swijsen, on the Segre manifold ($\mathbf{k}=(1,\ldots,1)$) with the Euclidean metric ($\alpha=1$) as part of his PhD thesis [77].
In the next subsection, we first determine a practical characterization of when two points in the pre-Segre–Veronese manifold can be connected with a geodesic. Then, in Section 4.2, we compute this minimizing geodesic if it exists. Finally, we compute minimizing geodesics in the Segre–Veronese manifold between two rank-$1$ tensors in Section 4.3 by appropriately lifting them to the pre-Segre–Veronese manifold. Additionally, an efficient matchmaking algorithm is presented to compute this lift, essentially solving a particular combinatorial problem in linear time.
4.1 Compatibility
Our main motivation for considering $\alpha$-warped geometries of the Segre–Veronese manifold is the observation in [77, Chapter 6] that not all pairs of points in the Segre manifold are connected by a minimizing geodesic. That is, the Segre manifold is not geodesically connected in the standard Euclidean geometry.
The main result of this subsection, Proposition 4.2, will show that the geodesic connectedness of $\mathcal{P}_{\alpha}^{\mathbf{k}}$ can be characterized by the following property.
Definition 4.1 ($\alpha$-compatibility).
Consider two points $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})$ in $\mathcal{P}_{\alpha}^{\mathbf{k}}$. Let
$M(p,q):=\mathrm{dist}_{\mathbb{S}^{\mathbf{n},\mathbf{k}}}((\mathbf{u}_{1},\ldots,\mathbf{u}_{d}),(\mathbf{v}_{1},\ldots,\mathbf{v}_{d}))=\sqrt{\sum_{i=1}^{d}k_{i}\cdot\sphericalangle^{2}(\mathbf{u}_{i},\mathbf{v}_{i})}.$
Then, $p$ and $q$ are called $\alpha$-compatible if $M(p,q)<\alpha^{-1}\pi.$
The dependency of $M(p,q)$ on the parameters $\mathbf{k}\geq 1$ will not be emphasized in the notation, instead being implicit in the fact that $p$, $q\in\mathcal{P}_{\alpha}^{\mathbf{k}}$.
Proposition 4.2.
Two points $p$ and $q$ can be connected by a minimizing geodesic in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ if and only if $p$ and $q$ are $\alpha$-compatible.
Proof 4.3.
Let $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$.
The case where $\mathbf{u}_{i}=\mathbf{v}_{i}$ for all $i=1,\dots,d$, so $M(p,q)=0$, corresponds to the case of straight line geodesics in Proposition 3.1. All such $p$ and $q$ can be connected by a straight line and since $M(p,q)=0$ they are always $\alpha$-compatible. Hence, the claim holds in this special case.
The remainder of the proof treats the case $M(p,q)>0$. For brevity, we let $M=M(p,q)$. Recall from [21, Lemma 4.4] that the projection of a geodesic in a warped product onto the second factor decreases in length if the projected curve in the second factor is replaced by a minimizing geodesic. This implies that minimizing geodesics in a warped product project to minimizing geodesics.
The geodesics passing through $p$ are described by Lemma 2.4. In particular, a minimizing geodesic $\sigma(s)=(\mu(s),\mathbf{v}_{1}(s),\dots,\mathbf{v}_{d}(s))$ that passes through both $p$ and $q$ must reach $q$ at $s=M$ because $\sigma$ is parameterized to unit spherical speed and the distance in $\mathbb{S}^{\mathbf{n},\mathbf{k}}$ between $(\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ and $(\mathbf{v}_{1},\ldots,\mathbf{v}_{d})$ is $M=M(p,q)$ from Definition 4.1. Plugging in this boundary condition $\sigma(M)=q$ allows us to determine the constant $B$ from Lemma 2.4:
$\frac{\mu(0)}{\mu(M)}=\frac{\lambda}{\mu}=\frac{\cos(B+\alpha M)}{\cos(B)}=\cos(\alpha M)-\sin(\alpha M)\tan(B),$
so that
(18) $B=\arctan\left(\cot(\alpha M)-\frac{\lambda\mu^{-1}}{\sin(\alpha M)}\right),$
provided that $\alpha M\neq\pi k$ for $k\in\mathbb{Z}$. Recall from the statement of Lemma 2.4 that if no such real $B$ exists, then there is also no geodesic $\sigma$ with unit spherical speed and $\sigma(0)=p$ and $\sigma(q)=M$. We can exploit this to prove the equivalence between compatibility and the existence of minimizing geodesics. We prove the two implications next, which will conclude the proof.
Geodesics $\Longrightarrow$ compatibility
Since $\sigma(s)$ exists by assumption, the solution $\mu(s)=\frac{\lambda\cos(B)}{\cos(\alpha s+B)}>0$ for all $0\leq s\leq M$. The inverse of the cosine has singularities at $B+\alpha s=\frac{\pi}{2}+k\pi$ for $k\in\mathbb{Z}$, so $\mu(s)$ cannot pass through them, and we must have $\frac{\pi}{2}+k\pi<B+\alpha s<\frac{\pi}{2}+k\pi$ for all $0\leq s\leq M$. In particular, plugging in the endpoint $s=0$ and combining with the bounds $\frac{\pi}{2}<B<\frac{\pi}{2}$ from Eq. 18 yields $k=0$. So, $\frac{\pi}{2}-B<\alpha s<\frac{\pi}{2}-B$ for all $0\leq s\leq M$. Plugging in the other endpoint $s=M$ and exploiting $\frac{\pi}{2}<B$, $M>0$, and $\alpha>0$ yields $0<\alpha M<\pi$. Hence, $p$ and $q$ are strictly $\alpha$-compatible.
Compatibility $\Longrightarrow$ geodesics
As $p$ and $q$ are $\alpha$-compatible, $0<\alpha M<\pi$, so $B$ is given as in Eq. 18. Since $\frac{\pi}{2}<B<\frac{\pi}{2}$, it follows from $0<\alpha M<\pi$ that
$B+\alpha M=\arctan\left(\cot(\alpha M)-\frac{\lambda\mu^{-1}}{\sin(\alpha M)}\right)+\alpha M<\arctan\left(\cot(\alpha M)\right)+\alpha M,$
because the arctangent is a monotonically increasing function and $\frac{\lambda\mu^{-1}}{\sin(\alpha M)}>0$ when $\alpha M\in(0,\pi)$. Exploiting that $\cot(\alpha M)=\tan(\frac{\pi}{2}-\alpha M)$, we get $\frac{\pi}{2}<B+\alpha M<\frac{\pi}{2}$ for all $0<\alpha M<\pi$. Consequently, $\frac{\pi}{2}<B+\alpha s<\frac{\pi}{2}$ for all $0\leq s\leq M$. Therefore, $\mu(s)$ is smooth curve that is strictly positive for all $0\leq s\leq M$. The curves $\mathbf{v}_{i}(s)$ are great circles, and we can always choose a parameterization so that $\mathbf{v}_{i}(0)=\mathbf{u}_{i}$ and $\mathbf{v}_{i}(M)=\mathbf{v}_{i}$. Hence, $\sigma\subset\mathcal{P}_{\alpha}^{\mathbf{k}}$ is a minimizing geodesic connecting $p$ and $q$.
The foregoing result implies that by changing the $\alpha$-warped metric of $\mathbb{R}_{>0}\times\mathbb{S}^{\mathbf{n},\mathbf{k}}$ we can modify whether two points $p$ and $q$ are connected by a minimizing geodesic. Specifically, we have the following result.
Proposition 4.4.
The pre-Segre–Veronese manifold $\mathcal{P}_{\alpha}^{\mathbf{k}}$ is geodesically connected if and only if $0<\alpha<\frac{1}{\sqrt{k_{1}+\cdots+k_{d}}}$.
Proof 4.5.
Let $p,q\in\mathcal{P}_{\alpha}^{\mathbf{k}}$. Then, $M^{2}(p,q)\leq\sum_{i=1}^{d}k_{i}\pi^{2}$, with equality attained at $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ and $q=(\lambda,-\mathbf{u}_{1},\ldots,-\mathbf{u}_{d})$. The result follows from Proposition 4.2.
4.2 Pre-Segre–Veronese manifolds
Having determined when two points in the pre-Segre–Veronese manifold $\mathcal{P}_{\alpha}^{\mathbf{k}}$ can be connected by a minimizing geodesic, we proceed by computing such a geodesic between two $\alpha$-compatible points.
Theorem 4.6 (Logarithmic map of $\mathcal{P}_{\alpha}^{\mathbf{k}}$).
Let
$p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}\quad\text{and}\quad q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$
be such that $\mathbf{u}_{i}\neq-\mathbf{v}_{i}$ for all $i=1,\dots,d$. Let $M=M(p,q)$ from Definition 4.1. If $p$ and $q$ are $\alpha$-compatible, then
$\log_{p}(q)=\bigl(\mu\cos(\alpha M)-\lambda,\dot{\mathbf{u}}_{1},\dots,\dot{\mathbf{u}}_{d}\bigr),$
where for $i=1,\dots,d$ we have
$\dot{\mathbf{u}}_{i}={\frac{\mu\sin(\alpha M)}{\lambda\alpha M}\log^{\mathbb{S}}_{\mathbf{u}_{i}}(\mathbf{v}_{i})}.$
In particular, if $M=0$, then $\log_{p}(q)=(\mu-\lambda,0,\dots,0).$
Proof 4.7.
The case $M=0$ follows immediately from Proposition 3.1. Hence, in the remainder of the proof we consider the case $M>0$.
Based on the form of the exponential map in Proposition 3.1, we conjecture that
$\log_{p}(q)=(\dot{\lambda},\dot{\mathbf{u}}_{1},\dots,\dot{\mathbf{u}}_{d})=(C,D\log^{\mathbb{S}}_{\mathbf{u}_{1}}(\mathbf{v}_{1}),\dots,D\log^{\mathbb{S}}_{\mathbf{u}_{d}}(\mathbf{v}_{d}))$
for some $C\in\mathbb{R}$ and $D>0$. Now we just have to solve $q={}\exp_{p}(\log_{p}q)$ for $C$ and $D$.
Recall the expression for $\exp_{p}$ from Proposition 3.1. The first component yields
(19) $(\lambda+C)^{2}+(\lambda\alpha DM)^{2}=\mu^{2},$
having used Definition 4.1 and $\|\dot{\mathbf{u}}_{i}\|=D\|\log^{\mathbb{S}}_{\mathbf{u}_{i}}(\mathbf{v}_{i})\|=D\sphericalangle(\mathbf{u}_{i},\mathbf{v}_{i})$. Note that $N$ from Proposition 3.1 equals $DM$. Since $M>0$, there exists at least one $1\leq i\leq d$ so that $\mathbf{u}_{i}\neq\mathbf{v}_{i}$. In all of these factors, we get from Proposition 3.1 that
$\mathbf{v}_{i}=\exp^{\mathbb{S}}_{\mathbf{u}_{i}}(a\dot{\mathbf{u}}_{i})=\exp^{\mathbb{S}}_{\mathbf{u}_{i}}(aD\log^{\mathbb{S}}_{\mathbf{u}_{i}}(\mathbf{v}_{i})),$
so that, applying $\log_{\mathbf{u}_{i}}$ on both sides, we obtain $aD=1$. Then, from the expression of $a$ in Proposition 3.1, we find
$\frac{1}{D}=a=\frac{1}{\alpha DM}\left(\frac{\pi}{2}-\arctan\left(\frac{\lambda+C}{\alpha\lambda DM}\right)\right).$
Clearing $D^{-1}$, multiplying by $\alpha M$ and taking cotangents, we find
(20) $\mathrm{cotan}(\alpha M)={}$ $\frac{\lambda+C}{\alpha\lambda DM}.$
The square system of polynomial equations Eqs. 19 and 20 in $C$ and $D$ together with the constraint $D>0$ defines the intersection of a circle with a line in the half-plane $D>0$. Ignoring the $D>0$ constraint, the two intersection points are
$C={}\varsigma\mu\cos(\alpha M)-\lambda\quad\text{and}\quad D={}\varsigma\frac{\mu\sin(\alpha M)}{\lambda\alpha M},\quad\text{for }\varsigma\in\{-1,1\}.$
By $\alpha$-compatibility we have $0<\alpha M<\pi$, and since we posited $D>0$, $\varsigma$ has to be $1$.
With the above values for $C$ and $D$, one verifies by direct computation that $\exp_{p}\circ\log_{p}$ is the identity, concluding the proof.
Remark 4.8.
We caution that if $\mathbf{u}_{i}=-\mathbf{v}_{i}$ for some $i$, then the logarithmic map is not well defined. However, $p$ and $q$ can still be connected by a geodesic. It is just that this geodesic is not unique. Choosing $\dot{\mathbf{u}}_{i}$ to be any vector orthogonal to $\mathbf{u}_{i}$ and with norm $\pi$ will produce such a geodesic.
An immediate consequence of the previous theorem is the following expression for the distance between points in the pre-Segre–Veronese manifold.
Corollary 4.9 (Distance between compatible points in $\mathcal{P}_{\alpha}^{\mathbf{k}}$).
Consider two points $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ that are $\alpha$-compatible. Let $M:=M(p,q)$ from Definition 4.1. Then,
(21) $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)={\|\log_{p}(q)\|_{p}}=\sqrt{\lambda^{2}-2\lambda\mu\cos(\alpha M)+\mu^{2}}.$
Proof 4.10.
By exploiting Eqs. 8, 4.6 and 4.1, this becomes a straightforward computation:
$\norm{\log_{p}(q)}_{p}^{2}$ $=(\mu\cos(\alpha M)-\lambda)^{2}+\alpha^{2}\lambda^{2}\sum_{i=1}^{d}k_{i}\frac{\mu^{2}\sin^{2}(\alpha M)}{\lambda^{2}\alpha^{2}M^{2}}\norm{\log_{\mathbf{u}_{i}}(\mathbf{v}_{i})}^{2}$
$=\mu^{2}\cos^{2}(\alpha M)-2\lambda\mu\cos(\alpha M)+\lambda^{2}+\frac{\mu^{2}\sin^{2}(\alpha M)}{M^{2}}\sum_{i=1}^{d}k_{i}\sphericalangle^{2}(\mathbf{u}_{i},\mathbf{v}_{i})$
$=\mu^{2}(\cos^{2}(\alpha M)+\sin^{2}(\alpha M))-2\lambda\mu\cos(\alpha M)+\lambda^{2},$
which concludes the proof.
A formula, equivalent to Eq. 21, is
(22) $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)=\sqrt{(\lambda-\mu)^{2}+4\lambda\mu\sin(\alpha M/2)},$
which is more numerically stable for points that are close together.
The distance between $p$ and $q$ in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ is thus given by the law of cosines applied to an imaginary triangle where one side has length $\lambda$, the other side has length $\mu$, and the angle is the spherical distance $M(p,q)$ multiplied by the warping factor $\alpha$. We thus immediately obtain the following result.
Corollary 4.11.
The distance $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)$ is monotonically increasing as a function of the spherical distance $M(p,q)$ for $0<\alpha M(p,q)<\pi$.
Interestingly, we can also determine the distance between incompatible points, even though minimizing geodesics do not exist between them. While a minimizing geodesic does not exist, there exists a limiting piecewise smooth curve that is the infimizer of the length functional in 2.5. In Fig. 1(b), for example, this limiting piecewise smooth curve is the straight line from $(1,0)$ to $(0,0)$, composed with the straight line from $(0,0)$ to $(0,1)$.
Proposition 4.12 (Distance between incompatible points in $\mathcal{P}_{\alpha}^{\mathbf{k}}$).
Two $\alpha$-incompatible points $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ are at
$\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)=\lambda+\mu.$
Proof 4.13.
We prove the distance is lower and upper bounded by $\lambda+\mu$.
Upper bound
The distance between $p$ and $q$, $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q):=\inf_{\gamma}\ell(\gamma),$ where $\gamma$ ranges over all piecewise smooth curves in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ with $\gamma(0)=p$ and $\gamma(1)=q$, can be upper bounded as follows. Let $\epsilon>0$ be arbitrary and consider the curve that goes first in a straight line from $p=(\lambda,\mathbf{u}_{1},\dots,\mathbf{u}_{d})$ to $(\epsilon,\mathbf{u}_{1},\dots,\mathbf{u}_{d})$, then continues on a circular arc to $(\epsilon,\mathbf{v}_{1},\dots,\mathbf{v}_{d})$, and finally continues in a straight line to $q=(\mu,\mathbf{v}_{1},\dots,\mathbf{v}_{d})$. It has length $\lambda+\mu+(\alpha M-2)\epsilon$, for all $\epsilon>0$. Hence, $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)\leq\lambda+\mu$, establishing the upper bound.
Lower bound
The set
$\mathcal{P}^{\prime}:=\{q^{\prime}\in\mathcal{P}_{\alpha}^{\mathbf{k}}\mid\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q^{\prime})<\lambda+\mu+1\}$
is an open subset of $\mathcal{P}_{\alpha}^{\mathbf{k}}$. Let $\gamma_{n}$ be a sequence of unit speed curves in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ that connect two incompatible points $p$ and $q$, and whose lengths approach the infimum. For sufficiently large $n$, we have $\gamma_{n}\subset\mathcal{P}^{\prime}$. Otherwise, $\gamma_{n}$ contains a point of $\mathcal{P}_{\alpha}^{\mathbf{k}}\setminus\mathcal{P}^{\prime}$, which lies at distance at least $\lambda+\mu+1$ from $p$, so $\ell(\gamma_{n})\geq\lambda+\mu+1\geq\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)+1$.
The limit $\gamma$ of the sequence of $\gamma_{n}$’s cannot be contained in $\mathcal{P}^{\prime}$, however, as this would show there is a piecewise smooth curve in $\mathcal{P}_{\alpha}^{\mathbf{k}}$ realizing the distance between $p$ and $q$, i.e., $p$ and $q$ would be $\alpha$-compatible by Proposition 4.2, contradicting their assumed incompatibility. As $\gamma_{n}\subset\mathcal{P}^{\prime}$, the limit $\gamma$ must lie in the (metric) closure
$\mathcal{C}=\mathcal{P^{\prime}}\cup\{0\}\cup\{q^{\prime}\in\mathcal{P}_{\alpha}^{\mathbf{k}}\mid\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q^{\prime})=\lambda+\mu+1\}$
of $\mathcal{P}^{\prime}$; 0 is in the closure since its distance to $p$ is at most $\lambda$. Since $\ell(\gamma)=\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)$, $\gamma\subset\mathcal{P}^{\prime}\cup\{0\}$, so $\gamma$ must contain 0. As $\gamma_{n}\to\gamma$, this entails that for a fixed $n$, there is a point $a=(\epsilon_{n},\mathbf{a}_{1},\dots,\mathbf{a}_{d})\in\mathcal{P}^{\prime}$ in $\gamma_{n}$ with $\epsilon_{n}$ minimal. Let $\gamma^{(1)}$ denote the part of $\gamma_{n}$ connecting $p$ and $a$ and let $\gamma^{(2)}$ be the part of $\gamma_{n}$ connecting $a$ and $q$. Then,
$\ell(\gamma_{n})=\ell(\gamma^{(1)})+\ell(\gamma^{(2)}).$
Moreover, if $\gamma^{(1)}(s)=(\nu(s),\mathbf{x}_{1}(s),\dots,\mathbf{x}_{d}(s))$, then
[…]
A similar argument shows that $\ell(\gamma^{(2)})\geq\mu-\epsilon_{n}$. Specifically, for every $n$, there exists an $\epsilon_{n}$ such that $\ell(\gamma_{n})\geq\lambda+\mu-2\epsilon_{n}$. Letting $n\to\infty$, we have $\epsilon_{n}\to 0$ because $\gamma$ contains 0, which shows that $\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q)\geq\lambda+\mu$.
4.3 Segre–Veronese manifolds
By leveraging Lemmas 2.3 and 4.6, minimizing geodesics in the Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ that connect two rank-$1$ tensors are readily obtained. Indeed, a minimizing geodesic connecting $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ is the pushforward of any geodesic whose length coincides with
$\min_{\begin{subarray}{c}p\in\otimes^{-1}(\mathpzc{P}),\,q\in\otimes^{-1}(\mathpzc{Q})\end{subarray}}\mathrm{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q).$
Corollary 4.11 then leads us naturally to the following terminology.
Definition 4.14 (Matched representatives).
Let $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$. We say that the representatives $p^{\star}\in\otimes^{-1}(\mathpzc{P})$ and $q^{\star}\in\otimes^{-1}(\mathpzc{Q})$ are matched if
$(p^{\star},q^{\star})\in\underset{\begin{subarray}{c}p\in\otimes^{-1}(\mathpzc{P}),\,q\in\otimes^{-1}(\mathpzc{Q})\end{subarray}}{\arg\min}M(p,q),$
where $M$ is the distance from Definition 4.1.
We are now ready to state and prove the main result about minimizing geodesics connecting two rank-$1$ tensors in the $\alpha$-warped Segre–Veronese manifold.
Theorem 4.15 (The logarithmic map of $\mathcal{S}_{\alpha}^{\mathbf{k}}$).
Let $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ be rank-$1$ tensors. If $p=(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})\in\otimes^{-1}(\mathpzc{P})$ and $q=(\mu,\mathbf{v}_{1},\ldots,\mathbf{v}_{d})\in\otimes^{-1}(\mathpzc{Q})$ are matched, $\alpha$-compatible representatives, then
$\log{P}\mathpzc{Q}=(\mathrm{d}_{p}\otimes)(\log_{p}q),$
where $\mathrm{d}_{p}\otimes:\mathrm{T}_{p}{\mathcal{P}_{\alpha}^{\mathbf{k}}}\to\mathrm{T}_{\mathpzc{P}}{\mathcal{S}_{\alpha}^{\mathbf{k}}}$ is the differential of $\otimes$.
Proof 4.16.
[…]
The following two results follow similarly.
Proposition 4.17.
Two rank-1 tensors $\mathpzc{P}$ and $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ can be connected by a minimizing geodesic in $\mathcal{S}_{\alpha}^{\mathbf{k}}$ if and only if $\mathpzc{P}$ and $\mathpzc{Q}$ have matched $\alpha$-compatible representatives.
Proposition 4.18 (Distance between points in $\mathcal{S}_{\alpha}^{\mathbf{k}}$).
Let $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$. If $p=(\lambda,\mathbf{u}_{1},\dots,\mathbf{u}_{d})\in\otimes^{-1}(\mathpzc{P})$ and $q=(\mu,\mathbf{v}_{1},\dots,\mathbf{v}_{d})\in\otimes^{-1}(\mathpzc{Q})$ are matched, $\alpha$-compatible representatives, then
$\operatorname{dist}_{\mathcal{S}_{\alpha}^{\mathbf{k}}}(\mathpzc{P},\mathpzc{Q})=\operatorname{dist}_{\mathcal{P}_{\alpha}^{\mathbf{k}}}(p,q).$
The straightforward approach to determine matched representatives of $\mathpzc{P}$, $\mathpzc{Q}\in\mathcal{S}_{\alpha}^{\mathbf{k}}$ consists of choosing an arbitrary $p^{\star}\in\otimes^{-1}(\mathpzc{P})$ and then computing all distances $M(p^{\star},q)$ for all $q\in\otimes^{-1}(\mathpzc{Q})$ and selecting a minimum. By Lemma 2.2, the possible $q$ can be labeled by $(\sigma_{1},\dots,\sigma_{d})\in\{\,-1,1\,\}^{d}$ such that $\sigma_{1}^{k_{1}}\cdots\sigma_{d}^{k_{d}}=1$. Thus, this strategy has a computational complexity of $\mathcal{O}(2^{d}\dim\mathcal{S}_{\alpha}^{\mathbf{k}})$, where $\mathcal{O}(\dim\mathcal{S}_{\alpha}^{\mathbf{k}})$ is the asymptotic cost for computing one distance $M(p^{\star},q)$. We now explain how a matched representative can be constructed efficiently in $\mathcal{O}(d\dim\mathcal{S}_{\alpha}^{\mathbf{k}})$ operations.
Let $q_{\mathrm{ref}}\in\otimes^{-1}(\mathcal{Q})$ be a reference preimage. We then observe that
$M(p^{\star},q)^{2}=M(p^{\star},q_{\mathrm{ref}})^{2}+\sum_{i=1}^{d}\frac{1-\sigma_{i}}{2}\Delta_{i}$
with $\Delta_{i}=k_{i}((\sphericalangle_{i}-\pi)^{2}-\sphericalangle_{i}^{2})$, where $\sphericalangle_{i}$ is the angle between the $i$th components of $p^{\star}$ and $q_{\mathrm{ref}}$. Minimizing $M(p^{\star},q)$ thus corresponds to maximizing $\sum_{i=1}^{d}{\sigma_{i}}\Delta_{i}$ over $\sigma_{i}\in\{-1,1\}$. An unconstrained optimum is
(23) $\sigma_{i}=\begin{cases}1,&\text{ if $\Delta_{i}$ is nonnegative},\\ -1,&\text{otherwise}.\end{cases}$
If this assignment also satisfies the constraint $\sigma_{1}^{k_{1}}\cdots\sigma_{d}^{k_{d}}=1$, so that $q\in\otimes^{-1}(\mathpzc{Q})$, then we are done. Otherwise, there is at least one smallest $|\absolutevalue{\Delta_{\ell}}$ among the odd $k_{\ell}$’s. The sign of the corresponding $\sigma_{\ell}$ should be flipped.
Proposition 4.19 (Matchmaking).
The algorithm described in the previous paragraph produces matched representatives.
Proof 4.20.
A detailed proof is given in Section A.3.
The construction of the matchmaking algorithm suggests that determining an analogue of Proposition 4.4 is more complicated for the Segre–Veronese manifold. Essentially, we have to consider the following problem: Given $p^{\star}\in\mathcal{P}_{\alpha}^{\mathbf{k}}$, what is the set of $q\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ such that the corresponding matched representatives $q^{\star}\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ are $\alpha$-compatible with $p^{\star}$, and when is this set the whole of $\mathcal{P}_{\alpha}^{\mathbf{k}}$? By our calculations, this question can be reduced to a combinatorial optimization problem. As we were unable to produce a closed form of its solution in all cases, we first present the following general result.
Proposition 4.21.
The Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ is geodesically connected if $0<\alpha<1/\sqrt{k_{1}+\dots+k_{d}}$ and is not geodesically connected if $2/\sqrt{k_{1}+\dots+k_{d}}\leq\alpha$.
Proof 4.22.
The first statement is a corollary of Lemmas 2.2 and 4.4.
To prove the second statement, assume $\alpha\geq 2/\sqrt{k_{1}+\dots+k_{d}}$ and pick $p$, $q\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ with angles $\sphericalangle_{1}=\dots=\sphericalangle_{d}=\frac{\pi}{2}$ between their components. These two representatives are matched. However, using $M(p,q)^{2}=(k_{1}+\dots+k_{d})\frac{\pi^{2}}{4}$ in Definition 4.1 shows that they are not $\alpha$-compatible.
For $1/\sqrt{k_{1}+\dots+k_{d}}<\alpha<2/\sqrt{k_{1}+\dots+k_{d}}$, geodesic connectedness is more complicated and depends also on the parity of the $k_{i}$’s. For example, we have the following basic special case.
Proposition 4.23.
Let $\mathbf{k}=(k_{1},\dots,k_{d})$ with $k_{i}$ even for all $i$. Then, the Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ is geodesically connected if and only if $0<\alpha<\frac{2}{\sqrt{k_{1}+\dots+k_{d}}}$.
Proof 4.24.
If the $k_{i}$’s are even, then $\sigma_{1}^{k_{1}}\cdots\sigma_{d}^{k_{d}}=1$ is automatically satisfied. The condition that two representatives $p$ and $q$ are matched is then equivalent to the requirement that the component angles satisfy $0\leq\sphericalangle_{i}\leq\frac{\pi}{2}$. The largest value that $M(p,q)^{2}$ can attain over all $p,q\in\mathcal{P}_{\alpha}^{\mathbf{k}}$ is then $(k_{1}+\dots+k_{d})\frac{\pi^{2}}{4}$, by choosing all angles equal to $\pi/2$ as in the proof of Proposition 4.21. Hence, all points of $\mathcal{P}_{\alpha}^{\mathbf{k}}$ are compatible if and only if $0<\alpha<2/\sqrt{k_{1}+\dots+k_{d}}$, concluding the proof.
The previous result shows that the lower bound on $\alpha$ in Proposition 4.21 can be sharp, while the upper bound on $\alpha$ in Proposition 4.21 can fail to be sharp. A full analysis is beyond the scope of this article.
5 Sectional curvature
Formulae for the sectional curvature of a manifold are useful in many situations, for example when approximating or compressing data that live in said manifold [42, 25, 83], or to ensure uniqueness of the Fréchet mean [45, 46, 3], as required in the experiment from Section 6. We compute these curvatures here for the $\alpha$-warped Segre–Veronese manifolds.
Consider first the $\alpha$-warped Segre manifold $\mathcal{S}_{\alpha}=\mathcal{S}_{\alpha}^{(1,\dots,1)}$ with corresponding pre-Segre manifold $\mathcal{P}_{\alpha}$. To compute its curvature, we recall the following general result about warped products.
Proposition 5.1 (O’Neill [68, 7.42]).
For $(a,b)\in\mathcal{M}\times_{f}\mathcal{N}$, let $\dot{\mathbf{x}}$, $\dot{\mathbf{y}}$, $\dot{\mathbf{z}}\in\mathrm{T}_{a}\mathcal{M}$ and $\dot{\mathbf{u}}$, $\dot{\mathbf{v}}$, $\dot{\mathbf{w}}\in\mathrm{T}_{b}\mathcal{N}$ be tangent vectors, which naturally lift to $\mathrm{T}_{(a,b)}(\mathcal{M}\times_{f}\mathcal{N})$. Let $R_{\mathcal{M}}$ and $R_{\mathcal{N}}$ be the Riemann curvature tensors on $\mathcal{M}$ and $\mathcal{N}$, respectively. Then, the Riemann curvature tensor $R$ on $\mathcal{M}\times_{f}\mathcal{N}$ is described by: [^1] Note that we use a definition of $R$ that is consistent with [52] rather than [68]. The result is that some signs differ in our expressions from theirs.
  1. 1.
    $R(\dot{\mathbf{x}},\dot{\mathbf{y}})\dot{\mathbf{z}}=R_{\mathcal{M}}(\dot{\mathbf{x}},\dot{\mathbf{y}})\dot{\mathbf{z}}$,
  2. 2.
    $R(\dot{\mathbf{u}},\dot{\mathbf{x}})\dot{\mathbf{y}}=-\frac{1}{f}\operatorname{Hess}[f](\dot{\mathbf{x}},\dot{\mathbf{y}})\dot{\mathbf{u}}$,
  3. 3.
    $R(\dot{\mathbf{x}},\dot{\mathbf{y}})\dot{\mathbf{u}}=0$,
  4. 4.
    $R(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{x}}=0$,
  5. 5.
    $R(\dot{\mathbf{x}},\dot{\mathbf{u}})\dot{\mathbf{v}}=-\frac{\left\langle{}\dot{\mathbf{u}},\dot{\mathbf{v}}\right\rangle}{f}\nabla_{\dot{\mathbf{x}}}(\operatorname{grad}f)$,
  6. 6.
    $R(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{w}}=R_{\mathcal{N}}(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{w}}+\frac{\norm{\operatorname{grad}f}^{2}}{f^{2}}\left(\left\langle{}\dot{\mathbf{u}},\dot{\mathbf{w}}\right\rangle\dot{\mathbf{v}}-\left\langle{}\dot{\mathbf{v}},\dot{\mathbf{w}}\right\rangle\dot{\mathbf{u}}\right)$.
Herein, $\left\langle{}\cdot,\cdot\right\rangle$ is the inner product on $\mathrm{T}_{(a,b)}(\mathcal{M}\times_{f}\mathcal{N})$, $\operatorname{grad}f$ is the Riemannian gradient of $f$, $\operatorname{Hess}[f]$ is the Riemannian Hessian of $f$, and $\nabla$ is the Levi–Civita connection.
For $\mathcal{S}_{\alpha}$, which locally looks like the warped product manifold $\mathcal{P}_{\alpha}$, these expressions simplify considerably. Applying Proposition 5.1 to $\mathcal{N}=\mathbb{S}^{\mathbf{n},\mathbf{1}}=\mathbb{S}^{n_{1}-1}\times\dots\times\mathbb{S}^{n_{d}-1}$ and $\mathcal{M}=\mathbb{R}_{>0}$ while observing that $\operatorname{grad}f=\alpha$ and $R_{\mathcal{M}}=0$ yields the next result.
Corollary 5.2.
In the notation of Proposition 5.1, the Riemann curvature tensor $R$ of $\mathcal{S}_{\alpha}$ at $\lambda\mathbf{u}_{1}\otimes\dots\otimes\mathbf{u}_{d}$ satisfies
$R(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{w}}=R_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}(\dot{\mathbf{u}},\dot{\mathbf{v}})\dot{\mathbf{w}}+\frac{1}{\lambda^{2}}(\left\langle{}\dot{\mathbf{u}},\dot{\mathbf{w}}\right\rangle\dot{\mathbf{v}}-\left\langle{}\dot{\mathbf{v}},\dot{\mathbf{w}}\right\rangle\dot{\mathbf{u}}),$
and it is 0 in the remaining directions.
As sectional curvatures can be defined in terms of the Riemann curvature tensor, we quickly find the next consequence of the last result.
Corollary 5.3.
For $(\mathbf{u}_{1},\dots,\mathbf{u}_{d})\in\mathbb{S}^{\mathbf{n},\mathbf{1}}$, let the lifts of the tangent vectors $\dot{\mathbf{u}}_{i}\in\mathrm{T}_{\mathbf{u}_{i}}\mathbb{S}^{n_{i}-1}$ and $\dot{\mathbf{v}}_{j}\in\mathrm{T}_{\mathbf{u}_{j}}\mathbb{S}^{n_{j}-1}$ be orthonormal with respect to $\langle\cdot,\cdot\rangle_{\mathcal{P}_{\alpha}}$. The sectional curvature $K$ of $\mathcal{S}_{\alpha}$ at $\lambda\mathbf{u}_{1}\otimes\dots\otimes\mathbf{u}_{d}$ satisfies
$K(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})=\begin{cases}\frac{1-\alpha^{2}}{\alpha^{2}\lambda^{2}},&\text{if $i=j$},\\ -\frac{1}{\lambda^{2}},&\text{otherwise},\end{cases}$
and $K=0$ in all remaining directions.
Proof 5.4.
Recall that $\left\langle{}\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j}\right\rangle_{\mathcal{P}_{\alpha}}=\alpha^{2}\lambda^{2}\left\langle{}\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j}\right\rangle_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}$. We compute
$K(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})={}$ $\left\langle{}R(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})\dot{\mathbf{v}}_{j},\dot{\mathbf{u}}_{i}\right\rangle_{\mathcal{P}_{\alpha}}$
$={}$ $\left\langle{}R_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})\dot{\mathbf{v}}_{j},\dot{\mathbf{u}}_{i}\right\rangle_{\mathcal{P}_{\alpha}}+\frac{1}{\lambda^{2}}\left(\left\langle{}\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j}\right\rangle_{\mathcal{P}_{\alpha}}^{2}-\left\langle{}\dot{\mathbf{v}}_{j},\dot{\mathbf{v}}_{j}\right\rangle_{\mathcal{P}_{\alpha}}\left\langle{}\dot{\mathbf{u}}_{i},\dot{\mathbf{u}}_{i}\right\rangle_{\mathcal{P}_{\alpha}}\right)$
$={}$ $\alpha^{2}\lambda^{2}\left\langle{}R_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}(\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{j})\dot{\mathbf{v}}_{j},\dot{\mathbf{u}}_{i}\right\rangle_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}-\frac{1}{\lambda^{2}}$
$={}$ $\frac{\alpha^{2}\lambda^{2}}{\alpha^{4}\lambda^{4}}\left\langle{}R_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}(\alpha\lambda\dot{\mathbf{u}}_{i},\alpha\lambda\dot{\mathbf{v}}_{j})\alpha\lambda\dot{\mathbf{v}}_{j},\alpha\lambda\dot{\mathbf{u}}_{i}\right\rangle_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}-\frac{1}{\lambda^{2}}$
$={}$ $\frac{1}{\alpha^{2}\lambda^{2}}K_{\mathbb{S}^{\mathbf{n},\mathbf{1}}}(\alpha\lambda\dot{\mathbf{u}}_{i},\alpha\lambda\dot{\mathbf{v}}_{j})-\frac{1}{\lambda^{2}}$
$={}$ $\frac{1}{\alpha^{2}\lambda^{2}}\delta_{ij}-\frac{1}{\lambda^{2}}.$
In all other directions the sectional curvature is 0, because $R=0$ in these directions.
The Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$, $\mathbf{k}=(k_{1},\dots,k_{d})$, is a Riemannian submanifold of the Segre manifold $\mathcal{S}_{\alpha}^{\mathbf{1}}$, where $\mathbf{1}$ is a vector of $k_{1}+\dots+k_{d}$ ones. If we look closely at the expressions in Propositions 3.1 and 3.4 for geodesics in $\mathcal{S}_{\alpha}^{\mathbf{1}}$, we see that starting out in a tangent direction to $\mathcal{S}_{\alpha}^{\mathbf{k}}\subset\mathcal{S}_{\alpha}^{\mathbf{1}}$ produces a path wholly contained in $\mathcal{S}_{\alpha}^{\mathbf{k}}$. We have thus proved the following result.
Proposition 5.5.
The Segre–Veronese manifold $\mathcal{S}_{\alpha}^{\mathbf{k}}$ is a totally geodesic submanifold of the Segre manifold $\mathcal{S}_{\alpha}^{\mathbf{1}}$, and so its sectional curvature is described by restricting $K$ in Corollary 5.3 to the appropriate tangent subspace.
6 Numerical experiment
We conclude with a numerical illustration of an example application that benefits from knowledge of the exponential and logarithmic maps of the $\alpha$-warped Segre manifold, derived in Corollaries 3.4 and 4.15. This application will also illustrate their computational efficacy.
The $\alpha$-warped geometries of Segre–Veronese manifolds $\mathcal{S}_{\alpha}^{\mathbf{k}}$ were implemented in Julia, including the exponential map, logarithmic map, and the geodesic distance. A modified version of our implementation of these geometries is available as part of the Manifolds.jl library [5]. The Julia source code for the experiments below is available publicly at https://gitlab.kuleuven.be/numa/public/alpha-warped-experiments.
Recall the consensus aggregation problem from Chafamo, Shanmugam, and Tokcan [16]: Given $M$ approximate tensor rank decompositions of the same tensor $\mathpzc{A}\in\mathbb{R}^{n_{1}\times\dots\times n_{d}}$, namely
$\mathpzc{A}\approx\mathpzc{A}^{(m)}:=\sum_{i=1}^{r}\mathpzc{A}_{i}^{(m)}:=\sum_{i=1}^{r}\mathbf{a}_{1i}^{(m)}\otimes\dots\otimes\mathbf{a}_{di}^{(m)},\quad m=1,\,\dots,\,$M,
estimate the true decomposition of $\mathpzc{A}$. A consensus aggregation algorithm was proposed in [16]; we give a simplified description of the essential ingredients of their method. While tensor rank decompositions usually have generically unique decompositions into a set of rank-$1$ tensors [60, 80, 6], to perform aggregation one needs to determine how these rank-$1$ tensors match up across multiple decompositions. For this reason, the rank-$1$ tensors $\mathpzc{A}_{i}^{(m)}$ are first matched using a ($k$-means) clustering algorithm in [16]. This strategy can work well if the decomposition problem is well conditioned, so that $\mathpzc{A}^{(m)}$ being close to $\mathpzc{A}$ will imply that the corresponding rank-$1$ tensors are close too [12]. After matching up the rank-$1$ tensors, the results are aggregated in [16] by setting
$\widehat{\mathbf{a}}_{i}^{l}={}$ $\operatorname{median}(\mathbf{a}^{(m)}_{li}:m=1,\,\dots,\,M),\quad i=1,\ldots,r,\;l=1,\ldots,d.$
The median of these vectors is taken elementwise. Note that it is possible for $\widehat{\mathbf{a}}_{i}^{l}$ to have different values if different representatives are chosen. This is circumvented in [16] by aggregating positive data, for which there is a natural, positive, representative. The aggregated tensor decomposition is then
$\mathpzc{A}\approx\widehat{\mathpzc{A}}:=\sum_{i=1}^{r}\widehat{\mathbf{a}}_{i}^{1}\otimes\dots\otimes\widehat{\mathbf{a}}_{i}^{d}.$
The method from [16] essentially takes a Euclidean midpoint of multiple points in spheres of different radii with the goal of averaging rank-$1$ tensors in the Segre manifold. From a geometric viewpoint, it is more natural to exploit the intrinsic geometry of the Segre manifold in its Euclidean metric, i.e., $\mathcal{S}_{1}^{(1,\dots,1)}$, rather than using the product Euclidean metric on the pre-Segre manifold $\mathcal{P}_{1}^{(1,\dots,1)}$. Recall from Lemma 2.2 that pre-Segre and Segre manifolds are locally isometric under the warped product metric on the pre-Segre manifold, but not under the standard product metric that [16] use. Therefore, we propose here to perform consensus aggregation by estimating each rank-$1$ tensor with the Fréchet mean on the warped Segre manifold $\mathcal{S}_{\alpha}=\mathcal{S}_{\alpha}^{(1,\dots,1)}$. Recall that the Fréchet mean or Riemannian center of mass of points $x_{1},\dots,x_{n}\in\mathcal{M}$ in a Riemannian manifold $\mathcal{M}$ is defined as
$\operatorname{mean}_{\mathcal{M}}(x_{1},\dots,x_{n})=\underset{p\in\mathcal{M}}{\arg\min}\sum_{i=1}^{n}\mathrm{dist}_{\mathcal{M}}(p,{x_{i}})^{2}.$
Its uniqueness has been investigated among others in [45, 46, 3]. For manifolds with negative sectional curvatures, the Fréchet mean is unique and the above defines a valid function. Hence, we propose to estimate $\mathpzc{A}$ instead as
[…]
To ensure that all points are compatible by Proposition 4.21, one can choose $\alpha<1/\sqrt{d}$. This guarantees that the Fréchet mean is well defined for all inputs. Note that significant warping may not be necessary if all inputs to the Fréchet mean are sufficiently close to one another in the angular distance of Definition 4.1.
We reproduce the first experiment from [16], using the authors’ code with the same parameter values and initialization seed. This synthetic experiment involves generating a reference rank-$9$ tensor rank decomposition of a $40\times 20\times 2000$ real-valued tensor $\mathpzc{A}$ with positive elements. Twenty noisy versions of $\mathpzc{A}$ are created by adding Poisson noise as described in [16]. They are then decomposed with the zero-inflated Poisson tensor factorization (ZIPTF) method from [16] from $10$ random starting points. We compare the following three methods of aggregating the results:
  1. S.
    No aggregation, i.e., the first ZIPTF decomposition of the first noisy tensor is used as “consensus”;
  2. C.
    C-ZIPTF is the consensus aggregation of all corresponding rank-$1$ tensors in the $10$ decompositions of one noisy tensor by the method of medians, as described previously and in [16]; and
  3. $\alpha$.
    F-ZIPTF operates like C-ZIPTF but aggregates the rank-$1$ tensors by our proposed method of Fréchet means. The latter is computed in the $\alpha$-warped geometry of the Segre manifold for various choices of $\alpha$.
The Fréchet mean is approximated by successive geodesic interpolation, as implemented in Manifolds.jl; see for example [19, 17, 18, 22]. To measure the accuracy of these methods, we measure the error between the reference tensor rank decomposition $\mathpzc{A}=\mathpzc{A}_{1}+\dots+\mathpzc{A}_{9}$ and an approximating decomposition $\widehat{\mathpzc{A}}=\widehat{\mathpzc{A}}_{1}+\dots+\widehat{\mathpzc{A}}_{9}\approx\mathpzc{A}$ as the mean of the relative Riemannian distances between the rank-$1$ tensors in both decompositions:
[…]
where $\norm{\cdot}_{F}$ is the Frobenius norm. The rank-$1$ tensors are first optimally matched before computing this error.
Figure 2: The mean relative distance from Eq. 24 between the reference tensor rank decomposition and the approximate decompositions computed from noisy samples of this tensor rank decomposition using a single ZIPTF [16], the consensus C-ZIPTF [16], and the Fréchet F-ZIPTF in several $\alpha$-warped geometries of the Segre manifold (this article) on the decomposition problem setup described in Section 6. In the figure’s horizontal axis, “S” refers to a single ZIPTF, “C” to C-ZIPTF, and the numerical values $\alpha=\frac{1}{\sqrt{5}},\dots,\frac{1}{\sqrt{2}},1,\sqrt{2},\dots,\sqrt{5},$ correspond to F-ZIPTF where the Fréchet mean was computed in the $\alpha$-warped geometry of the Segre manifold.
The results of our experiments are summarized in Fig. 2. We conclude that the accuracy of the consensus variants is better than a single ZIPTF. There does not appear to be a meaningful difference between C-ZIPTF and F-ZIPTF. Moreover, the experiments suggest that F-ZIPTF is quite robust to the choice of warping factor $\frac{1}{\sqrt{5}}\leq\alpha\leq\sqrt{5}$.
The exponential and logarithmic maps of the $\alpha$-warped Segre manifold can be computed very efficiently. In particular, to compute all Fréchet means of the groups of rank-$1$ tensors in all $9$ warped geometries required $21.7$ seconds and involved $14,400$ evaluations of both the exponential and logarithmic maps. On average, this is approximately $0.75$ milliseconds for each evaluation of either exponential or logarithmic maps. This experiment was performed on a computer system comprising an AMD Ryzen 7 5800X3D CPU (8 cores, up to 3.4 GHz clock speed, 96 MB L3 cache memory) and $4\times 32$ GB DDR4-3600 main memory, running Julia 1.12.4 on Ubuntu 22.04.3 LTS.
Acknowledgments
We thank André Uschmajew for editing our article and the two reviewers for their detailed feedback that improved this article, leading to further enhancements to the presentation of our results. We are especially grateful for the suggestions leading to clarifications in Section 6 and the inclusion of the left panel in Fig. 1, Lemma 2.4, a shorter proof of Corollary 4.9, Proposition 4.23, and an expanded version of Fig. 2. The suggestions of another anonymous reviewer, which led to clarifications about Riemannian geometry in Sections 2.1 and 2.2, a reformulation of our results in terms of exponential and logarithmic maps rather than geodesics in Propositions 3.1 and 4.6, and a shorter non-constructive proof of Theorem 4.6, are also acknowledged.
Appendix A Proof of the technical results
A.1 Proof of Lemma 2.2
Each $\imath_{\sigma}$ is an isometry because $\mathcal{P}_{\alpha}^{\mathbf{k}}$ is a product manifold where $\imath_{\sigma}$ acts only on the spheres and the maps $\mathrm{Id}$ and $\mathrm{Id}$ are basic isometries of the sphere with the standard Euclidean inner product, and hence also with any scaled inner product.
Normality, i.e., $\otimes(p)=\otimes(q)$ implies there exists a $\sigma$ such that $q=\imath_{\sigma}(p)$ [68, Appendix A], follows from the multilinearity of the tensor product $\otimes$ [33, Chapter 1].
By the same arguments as [7, Section 4.1], $\otimes$ is a smooth covering map. Hence, $\dim\mathcal{P}_{\alpha}^{\mathbf{k}}=\dim\mathcal{S}_{\alpha}^{\mathbf{k}}$ and the differential $\mathrm{d}\otimes:\mathrm{T}\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathrm{T}\mathcal{S}_{\alpha}^{\mathbf{k}}$ is everywhere left-invertible.
The key part is showing that $\otimes:\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathcal{S}_{\alpha}^{\mathbf{k}}$ is Riemannian, i.e., the metric of $\mathcal{S}_{\alpha}^{\mathbf{k}}$ is the pushforward of the one of $\mathcal{P}_{\alpha}^{\mathbf{k}}$, which is a straightforward computation. Let $\dot{\mathbf{u}}_{i}\in\mathrm{T}_{\mathbf{u}_{i}}\mathbb{S}^{n_{1}-1}=\mathbf{u}_{i}^{\perp}$ and $\dot{\lambda}\in\mathbb{R}$ be arbitrary. By differentiating $\otimes:\mathcal{P}_{\alpha}^{\mathbf{k}}\to\mathcal{S}_{\alpha}^{\mathbf{k}}$, we get
$(\mathrm{d}_{(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})}\otimes)(\dot{\lambda},\dot{\mathbf{u}}_{1},\ldots,\dot{\mathbf{u}}_{d})\\ =\dot{\lambda}\mathpzc{U}+\lambda(\dot{\nu}_{k_{1}}(\dot{\mathbf{u}}_{1})\otimes\mathbf{u}_{2}^{\otimes k_{2}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}})+\cdots+\lambda(\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d-1}^{\otimes k_{d-1}}\otimes\dot{\nu}_{k_{d}}(\dot{\mathbf{u}}_{d})),$
where $\mathpzc{U}$ and $\nu_{k}(\mathbf{u})$ are as in Eqs. 16 and 17. The $\alpha$-warped metric on $\mathcal{S}_{\alpha}^{\mathbf{k}}$ satisfies
$\varsigma^{\alpha}_{\lambda\mathpzc{U}}\left((\mathrm{d}\otimes)(\dot{x},\dot{\mathbf{u}}_{1},\ldots,\dot{\mathbf{u}}_{d}),(\mathrm{d}\otimes)(\dot{y},\dot{\mathbf{v}}_{1},\ldots,\dot{\mathbf{v}}_{d})\right)=\dot{x}\dot{y}+(\alpha\lambda)^{2}\langle\dot{\mathpzc{U}},\dot{\mathpzc{V}}\rangle,$
having dropped the subscript $(\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d})$ of the differential and where
$\dot{\mathpzc{U}}$ $=(\dot{\nu}_{k_{1}}(\dot{\mathbf{u}}_{1})\otimes\mathbf{u}_{2}^{\otimes k_{2}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}})+\cdots+(\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d-1}^{\otimes k_{d-1}}\otimes\dot{\nu}_{k_{d}}(\dot{\mathbf{u}}_{d})),$
$\dot{\mathpzc{V}}$ $=(\dot{\nu}_{k_{1}}(\dot{\mathbf{v}}_{1})\otimes\mathbf{u}_{2}^{\otimes k_{2}}\otimes\cdots\otimes\mathbf{u}_{d}^{\otimes k_{d}})+\cdots+(\mathbf{u}_{1}^{\otimes k_{1}}\otimes\cdots\otimes\mathbf{u}_{d-1}^{\otimes k_{d-1}}\otimes\dot{\nu}_{k_{d}}(\dot{\mathbf{v}}_{d})).$
Recall from [36] that the Euclidean inner product between rank-$1$ tensors satisfies $\langle\mathbf{u}_{1}\otimes\cdots\otimes\mathbf{u}_{d},\mathbf{v}_{1}\otimes\cdots\otimes\mathbf{v}_{d}\rangle=\prod_{i=1}^{d}\langle\mathbf{u}_{i},\mathbf{v}_{i}\rangle.$ Hence, rank-$1$ tensors are orthogonal to one another if they are orthogonal in at least one factor. We compute that $\langle\nu_{k_{i}}(\dot{\mathbf{u}}_{i}),\mathbf{u}_{i}^{\otimes k}\rangle=0,$ while $\langle\nu_{k_{i}}(\dot{\mathbf{u}}_{i}),\nu_{k_{i}}(\dot{\mathbf{v}}_{i})\rangle=k_{i}\langle\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{i}\rangle$ if $\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{i}\in\mathrm{T}_{\mathbf{u}_{i}}\mathbb{S}^{n}$. As a consequence, we can observe that the cross terms in $\langle\dot{\mathpzc{U}},\dot{\mathpzc{V}}\rangle$ vanish, so we are left with
$\langle\dot{\mathpzc{U}},\dot{\mathpzc{V}}\rangle=\sum_{i=1}^{d}\langle\dot{\nu}_{k_{i}}(\dot{\mathbf{u}}_{i}),\dot{\nu}_{k_{i}}(\dot{\mathbf{v}}_{i})\rangle=\sum_{i=1}^{d}k_{i}\langle\dot{\mathbf{u}}_{i},\dot{\mathbf{v}}_{i}\rangle.$
This shows that $\varsigma_{\lambda\mathpzc{U}}$ is the pushforward of $g_{\lambda,\mathbf{u}_{1},\ldots,\mathbf{u}_{d}}$ under $\otimes$, concluding the proof.
A.2 Proof of Lemma 2.3
Every smooth curve in $\mathcal{N}$ has such a unique smooth lift through a normal covering by [68, Lemma A.9]. Moreover, as $\phi$ is a local isometry that restricts to an isometry on evenly covered neighborhoods, the length of $\gamma$ and its unique lift $\widetilde{\gamma}$ in $\mathcal{M}$ coincide. Indeed, if we partition the lift $\widetilde{\gamma}$ into the curve segments $\widetilde{\gamma}_{i}$ such that $\widetilde{\gamma}_{i}\subset\mathcal{M}_{i}\subset\mathcal{M}$ with $\phi|_{\mathcal{M}_{i}}:\mathcal{M}_{i}\to\mathcal{N}_{i}$ an isometry (such a partition exists and is used to prove the existence of a unique lift in [68, Lemma A.9]), then
$\ell(\widetilde{\gamma})=\sum_{i=1}^{k}\ell(\widetilde{\gamma}_{i})=\sum_{i=1}^{k}\ell(\phi(\widetilde{\gamma}_{i}))=\ell(\gamma).$
This concludes the first part of the proof.
If there were a strictly shorter smooth curve $\widetilde{\gamma}^{\prime}\subset\mathcal{M}$ connecting the endpoints of the lift $\widetilde{\gamma}$, then by the same partitioning argument $\phi\circ\widetilde{\gamma}^{\prime}\subset\mathcal{N}$ would be a shorter piecewise smooth curve ($\phi$ is a local embedding) than $\gamma$, contradicting its minimality. This proves the second part.
A.3 Proof of Proposition 4.19
We want to maximize $f=\sum_{i=1}^{d}\sigma_{i}\Delta_{i}$ over $\sigma_{i}\in\set{-1,1}$ with the constraint that $\sigma_{1}^{k_{1}}\cdots\sigma_{d}^{k_{d}}=1$.
We can assume without loss of generality that all $k_{i}$’s are odd. Indeed, even $k_{i}$ do not impact the constraint, so we can choose the corresponding sign so that $|\Delta_{i}|=\sigma_{i}\Delta_{i}$, which is clearly optimal, also without constraint, because of the bilinear structure of $f$. The matching algorithm chooses these signs precisely in this way.
We can further assume without loss of generality that all $\Delta_{i}$ are nonzero, for otherwise we could take the unconstrained optimum Eq. 23 and simply swap the sign of a $\sigma_{j}$ for which $\Delta_{j}=0$ to obtain a feasible solution that is also globally optimal, which is what the matching algorithm does.
Let $(\varsigma_{1},\dots,\varsigma_{d})$ be a constrained optimizer of $f$ that differs in the least number of indices from the unconstrained optimizer $(\sigma_{1},\dots,\sigma_{d})$ from Eq. 23. Denote the number of different indices by $p=\sharp\{i\in\{1,\dots,d\}\mid\sigma_{i}\neq\varsigma_{i}\}$. Then, there are three cases:
  1. $p=0$.
    The unconstrained optimum given by Eq. 23 and chosen by the algorithm satisfies the constraint.
  2. $p=1$.
    If $f_{0}$ denotes the value of $f$ for the unconstrained optimum, such a constrained optimum has value $|f_{0}-2\absolutevalue{\Delta_{i}}$. Hence, $i$ must correspond to the minimal $|\absolutevalue{\Delta_{i}}$, which the matching algorithm indeed selects by construction.
  3. $p\geq 2$.
    Let $i\neq j$ be any two indices for which $\varsigma_{i}\neq\sigma_{i}$ and $\varsigma_{j}\neq\sigma_{j}$. By changing the signs of $\varsigma_{i}$ and $\varsigma_{j}$ to match those of $\sigma_{i}$ and $\sigma_{j}$, respectively, this increases the objective by $2|\Delta_{i}|+2|\Delta_{j}|$. Since $|\Delta_{i}|,|\Delta_{j}|>0$, this modified, feasible constrained point with p2p-2 different indices has a strictly higher objective value than $(\varsigma_{1},\dots,\varsigma_{d})$, which was supposed to be the constrained optimizer with minimal $p$. This is a contradiction, so this case cannot occur.
This concludes the proof.
A.4 Proof of Eq. 14
Let $C=\frac{\dot{\lambda}}{\alpha\lambda N}$, and recall that $B=\mathrm{arctan}(C)$ and $A=\lambda\cos(B)=\frac{\lambda}{\sqrt{1+C^{2}}}.$ Define
(25) $z(t):=\frac{\lambda+\dot{\lambda}t}{\alpha\lambda Nt}=\frac{1}{\alpha Nt}+C,\quad\text{so that}\quad z^{\prime}(t)=-\frac{1}{\alpha Nt^{2}}.$
In this notation, Eq. 14 becomes $\alpha s(t)=\frac{\pi}{2}-\arctan(z(t)),$ so that
$\alpha s^{\prime}(t)=-\frac{z^{\prime}(t)}{1+z^{2}(t)}.$
Since $\sqrt{\dot{\lambda}^{2}+(\alpha\lambda N)^{2}}=\alpha\lambda N\sqrt{1+C^{2}}$, Eq. 13 simplifies as follows:
$\alpha s^{\prime}(t)$ $=\cos^{2}(\alpha s(t)+B)A^{-1}\sqrt{\dot{\lambda}^{2}+(\alpha\lambda N)^{2}}$
$=\cos^{2}(\alpha s(t)+B)\alpha N(1+C^{2})$
$=\alpha N(1+C^{2})\cos^{2}\Bigl(\frac{\pi}{2}-(\arctan(z(t))-\arctan(C))\Bigr)$
$=\alpha N(1+C^{2})\sin^{2}(\arctan(z(t))-\arctan(C)).$
Using basic trigonometric identities, it is easy to obtain the next identity:
(26) $\sin(\arctan(x)-\arctan(y))=\frac{x-y}{\sqrt{1+x^{2}}\sqrt{1+y^{2}}}.$
This allows us to conclude the proof because
$\alpha s^{\prime}(t)=\alpha N(1+C^{2})\frac{(z(t)-C)^{2}}{(1+z^{2}(t))(1+C^{2})}=\alpha N\frac{(\alpha Nt)^{-2}}{1+z^{2}(t)}=-\frac{z^{\prime}(t)}{1+z^{2}(t)}.$
A.5 Proof of Eq. 15
Using the same notation as in Section A.4, we get that $\alpha s(1)=\frac{\pi}{2}-\arctan(z(1)),$ so that
$\gamma(1)=\mu(s(1))=\frac{\lambda\cos(B)}{\cos(\alpha s(1)+B)}$ $=\frac{\lambda}{\sqrt{1+C^{2}}}\cdot\frac{1}{\sin(\arctan(z(1))-\arctan(C))}$
$=\frac{\lambda}{\sqrt{1+C^{2}}}\frac{\sqrt{1+C^{2}}\sqrt{1+z^{2}(1)}}{z(1)-C}$
$=\alpha\lambda N\sqrt{1+z^{2}(1)},$
where we used Eq. 26 in the penultimate and Eq. 25 in the final equality. Since $z^{2}(1)=(\lambda+\dot{\lambda})^{2}(\alpha\lambda N)^{-2}$, the proof is concluded.